### v1

# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

Abstract
This paper studies the problem of recovering the hidden vertex correspondence between two correlated random graphs. We propose the partially correlated Erdős-Rényi graphs model, wherein a pair of induced subgraphs with a certain number are correlated. We investigate the information-theoretic thresholds for recovering the latent correlated subgraphs and the hidden vertex correspondence. We prove that there exists an optimal rate for partial recovery for the number of correlated nodes, above which one can correctly match a fraction of vertices and below which correctly matching any positive fraction is impossible, and we also derive an optimal rate for exact recovery. In the proof of possibility results, we propose correlated functional digraphs, which partition the edges of the intersection graph into two types of components, and bound the error probability by lower-order cumulant generating functions. The proof of impossibility results build upon the generalized Fano’s inequality and the recovery thresholds settled in correlated Erdős-Rényi graphs model.
Keywords— Graph alignments, information-theoretic thresholds, Erdős-Rényi random graphs, partial recovery, exact recovery
1 Introduction
Recently, there has been a surge in interest in the problems of detecting graph correlations and the alignments of two correlated graphs. These questions have emerged across various domains. For instance, in social networks, determining the similarity between friendship networks across different platforms has garnered attention [NS08, NS09].
In the realm of computer vision, where 3-D shapes are often represented as graphs with adjacency matrices, the identification of whether two graphs represent the same object holds significant importance in pattern recognition and image processing [BBM05, CSS06]. In computational biology, the representation of biological networks as graphs aids in understanding and quantifying their correlation [SXB08, VCL+11]. Furthermore, in natural language processing, the ontology alignment problem involves representing each sentence as a graph, with nodes denoting words. The task of determining whether a given sentence can be inferred from the text directly relates to graph matching problems [HNM05]. Numerous graph models exist, with the Erdős-Rényi random graph model being a prominent example, as proposed by [PA59] and [Gil59]:
Definition 1 (Erdős-Rényi graph).
The Erdős-Rényi random graph is the graph on $n$ vertices where each edge connects with probability $10<p<$ 1 independently. Let $\mathcal{G}(n,p)$ denote the distribution of Erdős-Rényi random graphs with $n$ vertices and edge connecting probability $p$.
While there are inherent disparities between the Erdős-Rényi random graph model and networks derived from real-world scenarios, comprehensively understanding the Erdős-Rényi graphs remains profoundly significant. This understanding serves as a pivotal step in transitioning from solving detection and matching problems on Erdős-Rényi graphs to addressing challenges inherent in practical applications. The graph alignment problem entails identifying latent vertex correspondences between two graphs based on their structures. Following [PG11], for two random graphs $G_{1},G_{2}$ with vertex sets $V(G_{1}),V(G_{2})$ and edge sets $E(G_{1}),E(G_{2})$, a typical correlated graph model is correlated Erdős-Rényi random graph model:
Definition 2 (Correlated Erdős-Rényi graphs).
Let $\pi$ denote a latent bijective mapping from $V(G_{1})$ to $V(G_{2})$. We say a pair of graphs $(G_{1},G_{2})$ are correlated Erdős-Rényi graphs if both marginal distributions are $\mathcal{G}(n,p)$ and each pair of edges $(uv,\pi(u)\pi(v))$ for $u,v\in V(G_{1})$ follows the correlated bivariate Bernoulli distribution with correlation coefficient $\rho$.
Given observations on $G_{1}$ and $G_{2}$ under the correlated Erdős-Rényi graphs model, the goal is to recover the latent vertex mapping $\pi$. To quantify the performance of an estimator $\hat{\pi}$, we consider the following two recovery criterion:
  • Partial recovery: given a constant $\delta\in(0,1)$, we say $\hat{\pi}$ succeeds for partial recovery if
    |$|\{v\in V(G_{1}):\pi(v)=\hat{\pi}(v)\}|\geq\delta|V(G_{1})|.$ (1)
  • Exact recovery: we say $\hat{\pi}$ succeeds for exact recovery if
    $\pi(v)=\hat{\pi}(v),\quad\forall~{}v\in V(G_{1}).$ (2)
The information-theoretic thresholds for partial and exact recoveries of $\pi$ between two correlated Erdős-Rényi graphs have been extensively studied in the recent literature.
  • Partial Recovery. [GML21] presented an impossibility result for partial recovery in the sparse regime characterized by constant average degree and correlation. [HM23] showed that $np(p\vee\rho)\gtrsim\log\left(1+\frac{\rho}{p}\right)\vee 1$ suffices for partial recovery, while $n\gtrsim d(p+\rho-p\rho\|p)\log n$ is necessary, where $d(p\|q)$ denotes the Kullback–Leibler (KL) divergence between Bernoulli distributions with mean $p$ and $q$, respectively. The recent work [WXY22] settled the sharp threshold for dense graphs with $\frac{p}{p\vee\rho}=n^{-o(1)}$ and the thresholds within a constant factor for sparse ones with $\frac{p}{p\vee\rho}=n^{-\Omega(1)}$. For the sparse case, [DD23b] proved a sharp threshold when $\frac{p}{p\vee\rho}=n^{-\alpha+o(1)}$ for $\alpha\in(0,1]$.
  • Exact Recovery. Based on the properties of the intersection graph under a permutation $\pi$, [CK16, CK17] showed that the Maximal Likelihood Estimator (MLE) achieves exact recovery and established an information-theoretical lower bound with a gap of $\omega(1)$. The results are sharpened by [WXY22] where the sharp threshold for exact recovery are derived.
While numerous studies have extensively investigated recovery procedures within the correlated Erdős-Rényi graphs model, it is however imperative to recognize that the signal present in many graph structures from realistic models is often inferior to that within the correlated Erdős-Rényi graph. This discrepancy emerges as many nodes in realistic graphs do not have corresponding nodes in the second correlated graphs. To offer a resolution to this concern, we propose the following model where on part of the nodes from two graphs are correlated.
Definition 3 (Partially correlated Erdős-Rényi graphs).
Let $S^{*}\subseteq V(G_{1})$ be a latent subset of vertices and $\pi^{*}$ be a latent injective mapping from $S^{*}$ to $V(G_{2})$. We say a pair of graphs $(G_{1},G_{2})$ are partially correlated Erdős-Rényi graphs if both marginal distributions are $\mathcal{G}(n,p)$ and each pair of edges $(uv,\pi^{*}(u)\pi^{*}(v))$ for $u,v\in S^{*}$ follows the correlated bivariate Bernoulli distribution with correlation coefficient $\rho$.
The case $S^{*}=V(G_{1})$ reduces to a pair of correlated Erdős-Rényi graphs in Definition 2. Under the model in Definition 3, given $S^{*}\subseteq V(G_{1})$ and the range of $\pi^{*}$ denoted by $T^{*}\subseteq V(G_{2})$, the induced subgraphs $G_{1}[S^{*}]$ and $G_{2}[T^{*}]$ are correlated Erdős-Rényi graphs on $m$ vertices. Therefore, the model can be equivalently constructed by planting correlated Erdős-Rényi graphs over a pair of independent Erdős-Rényi graphs.
In this paper, we investigate the information-theoretic thresholds for recovering the correlated nodes $S^{*}$ and the mapping $\pi^{*}$. For notational simplicity, we also refer to the problem as recovering $\pi^{*}$ while keeping $S^{*}$ implicit as the domain of $\pi^{*}$. The success criterion is similar to (1) and (2), where $V(G_{1})$ shall be replaced by $S^{*}$. However, due to the potential inconsistency between the domain of $\pi^{*}$ and the estimator $\hat{\pi}:\hat{S}\mapsto V(G_{2})$, we define their overlap by:
$\mathrm{overlap}(\pi^{*},\hat{\pi})\triangleq\frac{|v\in S^{*}\cap\hat{S}:\pi^{*}(v)=\hat{\pi}(v)|}{|S^{*}|}.$ (3)
With the notion of overlap, the success criterion is given by
  • Partial recovery: $\hat{\pi}$ succeeds if $\mathrm{overlap}(\pi^{*},\hat{\pi})\geq\delta$ for a given constant $\delta\in(0,1)$;
  • Exact recovery: $\hat{\pi}$ succeeds if $\mathrm{overlap}(\pi^{*},\hat{\pi})=1$.
1.1 Main Results
In this subsection, we present the main results of the paper. We first introduce some notations for the presentation of main theorems. Throughout the paper, we assume $0<\rho\leq 1$, $0<p\leq\frac{1}{2}$, and the cardinality $|S^{*}|=m$ is known. We further assume $p\geq\frac{1}{n}$ since otherwise partial recovery is impossible by [WXY22]. For a pair of Bernoulli random variables with means $p_{1},p_{2}$ and correlation $\rho$, their bivariate distribution is denoted as $\mathrm{Bern}(p_{1},p_{2},\rho)$. In our model, a pair of correlated edges $(e,\pi^{*}(e))\sim\mathrm{Bern}(p,p,\rho)$. Define $p_{ij}\triangleq\mathbb{P}\left[e=i,\pi^{*}(e)=j\right]$ for $i,j\in\{0,1\}$. Then
$p_{11}=p^{2}+\rho p(1-p),\quad p_{10}=p_{01}=(1-\rho)p(1-p),\quad p_{00}=(1-p)^{2}+\rho p(1-p).$
For a pair $(e,\pi^{*}(e))$ both edges present with probability $p_{11}$, while for $\pi(e)\neq\pi^{*}(e)$ both $e$ and $\pi(e)$ present with probability $p^{2}$. The relative signal strength present in correlated edges is denoted by $\gamma\triangleq\frac{p_{11}}{p^{2}}-1=\frac{\rho(1-p)}{p}$. It turns out that such reparametrization of the correlation coefficient is crucial in determining the fundamental limits of the graph alignment problem.
Let $\mathcal{S}_{n,m}$ denote the set of injective mappings $\pi:S\subseteq V(G_{1})\mapsto V(G_{2})$ with $|S|=m$. Our goal is to identify the minimum number of correlated nodes $m$ such that recovery of $\pi^{*}$ is possible. For the possibility results, we consider the estimator defined as
$\hat{\pi}=\mathop{\text{argmax}}\limits_{\pi\in\mathcal{S}_{n,m}}\sum_{u\neq v}\mathbbm{1}_{uv\in E(G_{1})}\mathbbm{1}_{\pi(u)\pi(v)\in E(G_{2})}.$ (4)
Next, we introduce our main theorems. Define $\phi(\gamma)\triangleq(1+\gamma)\log(1+\gamma)-\gamma$.
Theorem 1 (Partial recovery).
For any constant $\delta\in(0,1)$, there exists constant $c_{1}(\delta)$ such that, when $m\geq\frac{c_{1}(\delta)\log n}{p^{2}\phi(\gamma)}$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (4) satisfies
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})\geq\delta\right]=1-o(1).$
Furthermore, for any $c\in(0,1)$, there exists $c_{2}(c,\delta)$ such that, when $m\leq\frac{c_{2}(c,\delta)\log n}{p^{2}\phi(\gamma)}$, for any estimator $\hat{\pi}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-c,$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
The possibility result is presented in the minimax sense, while the impossibility result is under a Bayesian model. Hence, the threshold holds for both minimax and Bayesian risks. Theorem 1 implies, for the purpose of partial recovery, the threshold for the number of correlated nodes $m$ is of the order $\frac{\log n}{p^{2}\phi(\gamma)}$, beyond which partial recovery is possible and below which partial recovery is impossible. The dependency on the ambient graph order is only logarithmic, while the scale in terms of $p$ and $\rho$ is characterized by $\frac{1}{p^{2}\phi(\gamma)}$.
Theorem 2 (Exact recovery).
When $m\geq C\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\right)$, where $C$ is a universal constant, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (4) satisfies
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})=1\right]=1-o(1).$
Furthermore, for any $c\in(0,1)$, there exists a constant $c_{3}$ only depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\right)$, for any estimator $\hat{\pi}$
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<1\right]\geq 1-c,$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
Theorem 2 implies, for the purpose of exact recovery, the threshold for the number of correlated nodes $m$ is of the order $\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$. Under the weak signal regime $\gamma=O(1)$, we obtain the same rate as for partial recovery described in Theorem 1. Although the $\log n$ scaling has been observed in many other problems on random graphs, under the strong signal regime $\gamma=\omega(1)$, Theorem 2 highlights a transition from $\frac{\log n}{p^{2}\phi(\gamma)}$ to $\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$ if $\log^{2}\frac{1}{p}-\log^{2}\frac{1}{\rho}\gtrsim\log n$. In the latter regime, the difficulty is essentially the recovery of mapping given the sets of correlated nodes $(S^{*},T^{*})$. See more discussions in Section 4.
In comparison to prior work, our results of partial recovery in Theorem 1 match the thresholds established in [WXY22] up to a constant factor in both dense and sparse regimes for the special case $S^{*}=V(G_{1})$. Furthermore, the threshold $\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$ for exact recovery is derived from addressing the alignment problem for the subgraphs with the additional information on the domain and range of $\pi^{*}$, which applies the result in [WXY22].
1.2 Related Work
Graph sampling.
Graph sampling methodologies are often propelled by many practical factors. Most notably, these encompass data scarcity, high data acquisition costs [SWM05], and limited surveys of hidden structures [LF09, YML13, FH16]. In scenarios where observations are sampled from two large networks, it becomes unrealistic to presume that correlation exists among all nodes within the sampled subgraphs. As a result, a pair of partially correlated graphs emerge naturally. While the precise number of correlated nodes may not be accessible, we often have some partial knowledge on the scale. For instance, when the observations are induced subgraphs of randomly selected nodes, the number of correlated nodes follows a hypergeometric distribution that concentrates around the mean value.
Besides the recent literature on the graph alignment problem, the correlation detection is another related topic. Given a pair of graphs, their correlation detection is formulated as a hypothesis testing problem, wherein the null hypothesis assumes independent random graphs, while the alternative assumes edge correlation under a latent permutation. [BCL+19] proposed a hypothesis testing model for correlated Erdős-Rényi graphs and provided a pseudo-polynomial time algorithm for detection under certain conditions on the edge connection probability and average degree. [WXY23] established the sharp threshold for dense Erdős-Rényi graphs and determined the threshold within a constant factor for sparse Erdős-Rényi graphs. [DD23a] derived the sharp threshold for sparse Erdős-Rényi graphs by analyzing the densest subgraph. Additionally, [MWXY21] proposed a polynomial time algorithm for detection by counting trees when the correlation coefficient exceeds a constant value. It is natural to ask whether the correlation can be detected when only a subsample from the graphs is collected. The probabilistic model is similar to the one present in the current paper, and we leave the exploration as our future work.
Efficient algorithms and computational hardness.
Numerous algorithms have been developed for the recovery problem. For example, [YG13] analyzed the percolation graph matching algorithm, [BCL+19] analyzed the problem using subgraph matching techniques, and [MX20] obtained an algorithm for the seeded setting based on a delicate analysis of local neighborhoods. However, these algorithms may be computationally inefficient. There are several polynomial-time algorithms for recovery, catering to different regimes correlation coefficients $\rho$. These include works by [BES80, Bol82, DCKG19, GM20, DMWX21, MRT23, MWXY23, DL23, MS24]. For instance, [MWXY23] proposed a polynomial-time algorithm for recovery by counting chandeliers when the correlation coefficient $\rho>\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant introduced in [Ott48]. Additionally, [DL23] introduced an efficient iterative polynomial-time algorithm for sparse Erdős-Rényi graphs when the correlation coefficient is a constant.
It is postulated in [HS17, Hop18, KWB19] that the framework of low-degree polynomial algorithms effectively demonstrates computation hardness of detecting and recovering latent structures, and it bears similarities to sum-of-square methods [HKP+17, Hop18]. Based on the conjecture on the hardness of low-degree polynomial algorithms, [MWXY21] proved that there is no polynomial-time test or matching algorithm when the correlation coefficient satisfies $\rho^{2}\leq\frac{1}{\text{polylog}(n)}$. Furthermore, [DDL23] showed computation hardness for detection and exact recovery when $p=n^{-1+o(1)}$ and the correlation coefficient $\rho<\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant, suggesting that several polynomial algorithms may be essentially optimal.
The maximal overlap in the form of (4) is a test statistic which aims to identify the mapping that maximizes the edge correlation between two graphs. It is known that finding the maximal overlap is an instance of quadratic assignment problem (QAP) [PRW94], which is NP-hard to solve or to approximate [MMS10]. There are many studies aiming to detect or recover latent structures based on the maximal overlap statistics [CK16, CK17, BCL+19, MX20, DMWX21, WXY22, WXY23, HM23]. Finally, we mention that the recent work [DDG24] approximated the maximal overlap within a constant factor in polynomial-time for sparse Erdős-Rényi graphs, and [DGH23] established a sharp transition on approximating problem on the performance of online algorithms for dense Erdős-Rényi graphs.
Other graph models.
Many properties of the correlated Erdős-Rényi graphs model have been extensively investigated. However, the strong symmetry and tree-like structure inherent in this model distinguish it significantly from graph models encountered in practical applications. Therefore, it is crucial to explore more general graph models. One such model is inhomogeneous random graph model, where the edge connecting probability varies among edges in the graph [RS23, SPT23, DFW23]. Besides, geometric random graph model [WWXY22, BB23, SNL+23, GL24], planted cycle model [MWZ23, MWZ24], planted subhypergraph model [DMW23] and corrupt model [AH23] have also been subjects of recent studies.
1.3 Notations
For any $n\in\mathbb{N}$, let $[n]\triangleq\{1,2,...,n\}$. For any $a,b\in\mathbb{R}$, let $a\wedge b=\min\{a,b\}$ and $a\vee b=\max(a,b)$. We use standard asymptotic notation: for two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$, if $a_{n}\leq Cb_{n}$ for some absolute constant $C$ and for all $n$; $a_{n}=\Omega(b_{n})$ or $a_{n}\gtrsim b_{n}$, if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$ or $a_{n}\asymp b_{n}$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$.
For a given graph $G$, let $V(G)$ denote its vertex set and $E(G)$ denote its edge set. Let $\mathsf{{v}}(G)=|V(G)|$ denote the order of $G$ and $\mathsf{{e}}(G)=|E(G)|$ denote size of $G$. For a set $V$, let $\tbinom{V}{2}\triangleq\{\{x,y\}:x,y\in V,x\neq y\}$ denote the collection of all subsets of $V$ of cardinality two. We also write $uv$ to denote an edge $\{u,v\}$. The induce subgraph of $G$ over a vertex set $V$ is denoted by $G[V]$. Given an injective mapping of vertices $\pi:S\subseteq V(G_{1})\mapsto V(G_{2})$, the induced injective mapping of edges is defined as $\pi^{{\mathsf{E}}}:\binom{S}{2}\mapsto\binom{V(G_{2})}{2}$ as $\pi^{{\mathsf{E}}}(uv)=\pi(u)\pi(v)$ for $u,v\in S$. We also succinctly write $\pi(e)=\pi^{{\mathsf{E}}}(e)$ for an edge $e$ when the meaning is clear from the context.
2 Correlated functional digraph
A mapping from a set to itself can be graphically represented as functional digraph (see, e.g., [Wes21, Definition 1.3.3]). Here we extend the notion to a mapping with different domain and range sets, where the elements from the two sets are correlated. While our focus in this section is on the mapping between the edges in $G_{1}$ and $G_{2}$, the graphical representation can be easily extended to mappings between two arbitrary finite sets such as vertices.
We first provide an equivalent description of the estimator in (4). Given a domain subset $S\subseteq\binom{V(G_{1})}{2}$ and an injective function $\pi:S\mapsto\binom{V(G_{2})}{2}$, we define the intersection graph ${\mathcal{H}}_{\pi}$ as
$V({\mathcal{H}}_{\pi})=V(G_{1}),\quad e\in E({\mathcal{H}}_{\pi})\text{ if andonly if }e\in E(G_{1})\cap S\text{ and }\pi(e)\in E(G_{2}).$
The estimator (4) maximizes the size of the intersection graph $|E({\mathcal{H}}_{\pi})|$. More generally, in our analysis in Section 3, we need to count the number of edges present in some subset ${\mathcal{E}}\subseteq S$ given by
$|{\mathcal{E}}\cap E({\mathcal{H}}_{\pi})|=\sum_{e\in{\mathcal{E}}}{\mathbf{1}_{\left\{{e\in E({\mathcal{H}}_{\pi})}\right\}}}=\sum_{e\in{\mathcal{E}}}{\mathbf{1}_{\left\{{e\in E(G_{1})}\right\}}}{\mathbf{1}_{\left\{{\pi(e)\in E(G_{2})}\right\}}}.$ (5)
Due to the correlation between the edges in $G_{1}$ and $G_{2}$, the counters ${\mathbf{1}_{\left\{{e\in E({\mathcal{H}}_{\pi})}\right\}}}$ are correlated random variables. The main idea is to decompose ${\mathcal{E}}$ into independent parts. Specially, the correlation is prescribed by the underlying mapping $\pi^{*}$ as illustrated in Figure 1, where the correlated edges are red dashed lines. To formally describe all correlation relationships, we introduce the correlated functional digraph of a mapping $\pi$ between a pair of graphs.
[…]
Figure 1: Examples of the mapping $\pi$ and the underlying correlation $\pi^{*}$, where the domain and range of $\pi$ and $\pi^{*}$ could be different.
Definition 4 (Correlated functional digraph).
Let $\pi^{*}:S^{*}\mapsto T^{*}$ be the underlying mapping between correlated elements. The correlated functional digraph of the function $\pi:S\mapsto T$ is constructed as follows. Let the vertex set be $S\cup S^{*}\cup T\cup T^{*}$. We first add every edge $e\mapsto\pi(e)$ for $e\in S$, and then merge each pair of nodes $(e,\pi^{*}(e))$ for $e\in S^{*}$ into one node.
[…]
Figure 2: The connected components in the correlated functional digraph.
It should be noted that both $\pi$ and $\pi^{*}$ are injective mappings under our model. After merging all pairs of nodes under $\pi^{*}$, the degree of each vertex in the correlated functional digraph is at most two. Therefore, the connected components consist of paths and cycles, where the self-loop is understood as a cycle of length one. The connected components are illustrated in Figure 2. Let ${\mathcal{P}}$ and ${\mathcal{C}}$ denote the collections of subsets of ${\mathcal{E}}$ belonging to different connected paths and cycles, respectively. Note that the sets from ${\mathcal{P}}$ and ${\mathcal{C}}$ are disjoint. Consequently,
$|{\mathcal{E}}\cap E({\mathcal{H}}_{\pi})|=\sum_{P\in{\mathcal{P}}}|P\cap E({\mathcal{H}}_{\pi})|+\sum_{C\in{\mathcal{C}}}|C\cap E({\mathcal{H}}_{\pi})|,$
where the summands are mutually independent.
In our model, the edge correlations are assumed to be homogeneous, and hence the distribution of $|P\cap E({\mathcal{H}}_{\pi})|$ and $|C\cap E({\mathcal{H}}_{\pi})|$ only depends on the size of the component. Let $\kappa^{{\mathsf{P}}}_{\ell}(t)$ and $\kappa^{{\mathsf{C}}}_{\ell}(t)$ denote the cumulant generating functions of $|P\cap E({\mathcal{H}}_{\pi})|$ and $|C\cap E({\mathcal{H}}_{\pi})|$ with $|P|=|C|=\ell$, respectively, and we have
$\log\mathbb{E}\left[e^{t|P\cap E({\mathcal{H}}_{\pi})|}\right]=\kappa^{{\mathsf{P}}}_{|P|}(t),\qquad\log\mathbb{E}\left[e^{t|C\cap E({\mathcal{H}}_{\pi})|}\right]=\kappa^{{\mathsf{C}}}_{|C|}(t).$
The lower-order cumulants can be promptly calculated. For instance,
$\kappa^{{\mathsf{C}}}_{1}(t)$ $=\log(1+p_{11}(e^{t}-1)),$ (6)
$\kappa^{{\mathsf{C}}}_{2}(t)$ $=\log(1+2p^{2}(e^{t}-1)+p_{11}^{2}(e^{t}-1)^{2}).$ (7)
It is however essential to establish upper bounds for higher-order cumulants in terms of lower-order ones. To this end, we introduce the following lemma.
Lemma 1.
For any $\rho>0$, $10<p<$ 1, and $t>0$,
$\kappa^{{\mathsf{P}}}_{1}(t)\leq\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\leq\kappa^{{\mathsf{C}}}_{1}(t)\quad\text{and}\quad\kappa^{{\mathsf{P}}}_{\ell}(t)\leq\kappa^{{\mathsf{C}}}_{\ell}(t)\leq\frac{\ell}{2}\kappa^{{\mathsf{C}}}_{2}(t),\quad\forall~{}\ell\geq 2.$
Consequently,
$\log\mathbb{E}\left[e^{t|{\mathcal{E}}\cap E({\mathcal{H}}_{\pi})|}\right]\leq\frac{|{\mathcal{E}}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$ (8)
where $L$ denotes the number of self-loops.
The proof of Lemma 1 is deferred to Section A.1. The special case that both $\pi$ and $\pi^{*}$ are bijective has been studied in [WXY22, DD23b, HM23], the correlation relationships under which can be characterized by a permutation $(\pi^{*})^{-1}\circ\pi$. In this case, the connected components of the functional digraph of permutations are all cycles. However, in our case, the domain and range of $\pi$ and $\pi^{*}$ could be different and we need to deal with delicate correlations among the edges involving both cycles and paths by Lemma 1.
3 Recovery by maximizing the size of intersection graph
In this section, we prove the possibility results by analyzing the estimator $\hat{\pi}$ given in (4). By the optimality condition, it suffices to show that, for any $\pi^{*}\in\mathcal{S}_{n,m}$, we have $e({\mathcal{H}}_{\pi^{*}})$ exceeds $\max_{\pi:d(\pi,\pi^{*})>\tau}e({\mathcal{H}}_{\pi})$ with high probability when the underlying correlation is specified by $\pi^{*}$, where the thresholds $\tau=0$ and $\delta m$ are for exact and partial recoveries, respectively. In the following, we fix $\pi^{*}$ and provide a general recipe for the upper bound of $\mathbb{P}_{\pi^{*}}[d(\hat{\pi},\pi^{*})=k]$. The overall error probability follows from the summation over the desired range of $k$.
Let $\mathcal{T}_{k}\subseteq\mathcal{S}_{n,m}$ denote the set of injections $\pi$ such that $d(\pi,\pi^{*})=k$. For $\pi\in\mathcal{T}_{k}$, by definition, there exists a set of correctly matched vertices (the self-loops in the correlated functional digraph of $\pi$ over the vertices), denoted by $F_{\pi}\triangleq\{v\in S^{*}\cap S:\pi^{*}(v)=\pi(v)\}$ of cardinality $|F_{\pi}|=m-k$. The induced subgraphs of ${\mathcal{H}}_{\pi}$ and ${\mathcal{H}}_{\pi^{*}}$ over $F_{\pi}$ are identical. Therefore,
$e({\mathcal{H}}_{\pi})\geq e({\mathcal{H}}_{\pi^{*}})\iff e({\mathcal{H}}_{\pi})-e({\mathcal{H}}_{\pi}[F_{\pi}])\geq e({\mathcal{H}}_{\pi^{*}})-e({\mathcal{H}}_{\pi^{*}}[F_{\pi}]).$
It should be noted that correlated random variables are contained within the two sides of the inequality. Nevertheless, for any threshold $\tau_{k}$, either $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}$ or $\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}$ holds. Therefore, we have the following upper bound:
$\{d(\hat{\pi},\pi^{*})=k\}\subseteq\bigcup_{\pi\in{\mathcal{T}}_{k}}\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\}\cup\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\}.$
The first event is indicative of a weak signal, while the latter implies the presence of strong noise. The crucial result to establish is that, for a suitable threshold $\tau_{k}$, both bad events will occur with a low probability. Here we may pick $\tau_{k}$ a function of all other parameters $m,k,p,\rho$. For brevity we also write $\tau_{k}=\tau(m,k,p,\rho)$.
Bad event of signal.
For a fixed $\pi\in\mathcal{T}_{k}$, the random variable $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])$ counts the total number of edges among $N_{k}\triangleq\binom{m}{2}-\binom{m-k}{2}=mk(1-\frac{k+1}{2m})$ pairs of vertices, where each edge presents independently with probability $p_{11}$. Furthermore, $F_{\pi}$ is a subset of $S^{*}$ of cardinality $m-k$. While the size of ${\mathcal{T}}_{k}$ could be large, the total number of possible $F_{\pi}$ is at most $\binom{m}{m-k}=\binom{m}{k}$. Therefore,
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\mathbb{P}\left[\bigcup_{\begin{subarray}{c}F\subseteq S^{*}\\ |F|=m-k\end{subarray}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F])<\tau_{k}\right\}\right]$
$\leq\binom{m}{k}\mathbb{P}\left[\mathrm{Bin}(N_{k},p_{11})<\tau_{k}\right].$ (9)
For $\tau_{k}<N_{k}p_{11}$, the tail of binomial distributions follows from the standard Chernoff bound.
Bad event of noise.
The analyses for the noise part is more involved due to the mismatch between $\pi$ and the underlying $\pi^{*}$. Let $S_{\pi}$ denote the domain of $\pi$, and ${\mathcal{E}}_{\pi}\triangleq\binom{S_{\pi}}{2}-\binom{F_{\pi}}{2}$. Then the total number of edges $\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])$ can be equivalently represented as $|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|$, and the cumulant generating function has been upper bounded in Lemma 1 thanks to the decomposition based on the correlated functional digraph. Thus, the error probability can be obtained via the Chernoff bound by optimizing over $t>0$ in (8).
To this end, we need to upper bound the number of self-loops in (8). For a self-loop over an edge $e=uv$, we have $\pi(uv)=\pi^{*}(uv)$. Note that ${\mathcal{E}}_{\pi}$ excludes the edges in the induced subgraph over $F_{\pi}$. It necessarily holds that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$, which contributes two mismatched vertices in the reconstruction of the underlying mapping. Since the total number of mismatched vertices for $\pi\in{\mathcal{T}}_{k}$ equals to $k$, the number of self-loops is at most $\frac{k}{2}$. Consequently, applying (8) with the formula of lower-order cumulants (6) and (7) yields the following lemma, whose proof is deferred to Section A.2.
Lemma 2.
If $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$, then
$\mathbb{P}\left[|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|\geq\tau_{k}\right]\leq\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}+\frac{k\gamma}{4(2+\gamma)}\right).$
In view of Lemma 2, for $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$, we can apply the union bound for the probability of the bad event due to noise. It remains to upper bound the cardinality of $\mathcal{T}_{k}$. We first choose $m-k$ elements from the domain of $\pi^{*}$ and map them to the same value as $\pi^{*}$. Then, the remaining domain and range of size $k$ and the mapping are selected arbitrarily. Then we obtain
$|\mathcal{T}_{k}|\leq\tbinom{m}{m-k}\tbinom{n-m+k}{k}^{2}k!\leq\frac{m^{k}n^{2k}}{k!^{2}},$
where the last step applies the upper bound $\binom{n}{k}\leq\frac{n^{k}}{k!}$. Since $e^{\frac{k\gamma}{4(2+\gamma)}}\leq e^{k/4}$, we have
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$ $=\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|\geq\tau_{k}\right\}\right]$
$\leq|{\mathcal{T}}_{k}|e^{\frac{k\gamma}{4(2+\gamma)}}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right)$
$\leq n^{3k}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right).$ (10)
The following propositions provide sufficient conditions on $m$ for partial and exact recoveries.
Proposition 1 (Upper bound for partial recovery).
For any $\delta\in(0,1)$, there exists a constant $c_{1}(\delta)>0$ such that, when $m\geq\frac{c_{1}(\delta)\log n}{p^{2}\phi(\gamma)}$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (4) satisfies
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\leq(\log n)^{-1+o(1)}.$
Proposition 2 (Upper bound for exact recovery).
There exists a universal constant $C>0$ such that, when $m\geq C\left(\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\vee\frac{\log n}{p^{2}\phi(\gamma)}\right)$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (4) satisfies
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\leq\frac{\exp(-\log m)}{1-\exp(-\log m)}+\frac{\exp(-\log n)}{1-\exp(-\log n)}.$
By Propositions 1 and 2, we prove the possibility results in Theorems 1 and 2.
3.1 Proof of Proposition 1
Let $\tau_{k}=N_{k}p_{11}(1-\eta)$ with
$\eta=\sqrt{\frac{8h(\frac{k}{m})}{kp_{11}}}\cdot\mathbbm{1}_{k\leq m-1}+\sqrt{\frac{\log n}{kmp_{11}}}\cdot\mathbbm{1}_{k=m},$
where $h(x)\triangleq-x\log x-(1-x)\log(1-x)$ is the binary entropy function. Since $h(x)/x$ decreases in $(0,1)$, $h(x)/x\geq h(1-\delta)/(1-\delta)$ for $11-\delta\leq x<$ 1. By Lemma 17.5.1 in [CT06], we have $\tbinom{m}{k}\leq\exp\left[mh(k/m)\right]$ for any $k\leq m-1$. By (9) and the Chernoff bound (21), when $k\leq m-1$, we have
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]\leq\binom{m}{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right)$
$\leq$ $\exp\left(mh\left(\frac{k}{m}\right)-\frac{N_{k}p_{11}\eta^{2}}{2}\right)\leq\exp\left(-mh\left(\frac{k}{m}\right)\right).$
When $k=m$, since $N_{k}=\frac{mk}{2}\left(2-\frac{k+1}{m}\right)\geq\frac{mk}{3}$, we have
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]\leq\binom{m}{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right)\leq\exp\left(-\frac{\log n}{6}\right).$
Pick $c_{1}(\delta)=100\vee\frac{200h(1-\delta)}{1-\delta}$. We then verify the condition in Lemma 3: $\eta\leq\frac{\gamma}{4(1+\gamma)}$. Since $p^{2}\phi(\gamma)\leq p^{2}\gamma(1+\gamma)\leq 1$, we get $m\geq c_{1}(\delta)\log n$. Therefore,
$\eta$ $\leq\sqrt{\frac{8h(1-\delta)}{(1-\delta)mp_{11}}}\cdot\mathbbm{1}_{k\leq m-1}+\sqrt{\frac{\log n}{m^{2}p_{11}}}\cdot\mathbbm{1}_{k=m}\leq\left(\sqrt{\frac{8h(1-\delta)}{1-\delta}}\vee\frac{1}{\sqrt{c_{1}(\delta)}}\right)\frac{1}{\sqrt{mp_{11}}}$
$\leq\left(\sqrt{\frac{8h(1-\delta)}{(1-\delta)c_{1}(\delta)}}\vee\frac{1}{c_{1}(\delta)}\right)\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}}\leq\frac{1}{5}\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}}.$
Recall the assumption stated in Section 1.1, where it’s asserted that $p\geq n^{-1}$, thereby implying $\log(1+\gamma)\leq\log n$. When $\gamma>10$, $\eta\leq\frac{1}{5}\leq\frac{\gamma}{4(1+\gamma)}$. When $\gamma\leq 10$, since $\log(1+x)-\frac{x}{1+x}-x^{2}\leq 0$ for any $x>0$, $\sqrt{\frac{1}{mp_{11}}}\leq\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}}\leq\frac{\gamma}{\sqrt{\log n}}\leq\frac{\gamma}{4(1+\gamma)}$. Therefore, we obtain $\eta\leq\frac{\gamma}{4(1+\gamma)}$. By Lemma 3, $\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}=(1+\gamma)(1-\eta)>1$. Applying Lemma 2, we derive (10). Combining this with (19) in Lemma 3, we obtain
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $n^{3k}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right)$
$=$ $n^{3k}\exp\left\{-\frac{N_{k}p^{2}}{2}\phi\left[(1+\gamma)(1-\eta)-1\right]\right\}\leq n^{3k}\exp\left(-\frac{N_{k}p^{2}}{8}\phi(\gamma)\right).$
Sum over $k\geq(1-\delta)m$, since $N_{k}\geq\frac{km}{3}$, we obtain
$\sum_{k=\delta m}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]$
$\leq$ $\sum_{k=\delta m}^{m}\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]+\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $\exp\left(-\frac{\log n}{6}\right)+\sum_{k=(1-\delta)m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]+\sum_{k=(1-\delta)m}^{m}\left[n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\right]^{k}$
$\leq$ $\exp\left(-\frac{\log n}{6}\right)+\frac{\exp\left[-(1-\delta)m\log n\right]}{1-\exp(-\log n)}+\sum_{k=(1-\delta)m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right].$
Combining this with Lemma 4, $\sum_{k=(1-\delta)m}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]\leq(\log n)^{-1+o(1)}$. We finish the proof.
3.2 Proof of Proposition 2
Let $\tau_{k}=N_{k}p_{11}(1-\eta)$ with $\eta=\frac{\gamma}{4(1+\gamma)}$, by (9) and the Chernoff bound (21),
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]\leq\binom{m}{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right).$
By Lemma 3, $\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}=(1+\gamma)(1-\eta)>1$. Applying Lemma 2, we derive (10). Combining this with (19) in Lemma 3, we obtain
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $n^{3k}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right)$
$=$ $n^{3k}\exp\left\{-\frac{N_{k}p^{2}}{2}\phi\left[(1+\gamma)(1-\eta)-1\right]\right\}\leq n^{3k}\exp\left[-\frac{N_{k}p^{2}}{8}\phi(\gamma)\right].$
Sum over $k\geq 1$, since $N_{k}\geq\frac{km}{3}$ and $\tbinom{m}{k}\leq m^{k}$, we obtain
$\sum_{k=1}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]$
$\leq$ $\sum_{k=1}^{m}\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]+\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $\sum_{k=1}^{m}\left[m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\right]^{k}+\left[n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\right]^{k}.$
Pick the universal constant $C=400$. Recall that $\phi(\gamma)=(1+\gamma)\log(1+\gamma)-\gamma$. When $\gamma\leq 1$, since $\phi(\gamma)\leq\frac{\gamma^{2}}{1+\gamma}$, we obtain $p^{2}\phi(\gamma)\leq 16p_{11}\eta^{2}$. Therefore, $m\geq\frac{400\log n}{p^{2}\phi(\gamma)}$ implies $m\geq\frac{12\log m}{p_{11}\eta^{2}}$. When $\gamma>1$, since $\gamma\leq 32(1+\gamma)\left[\frac{\gamma}{4(1+\gamma)}\right]^{2}$, we obtain $p^{2}\gamma\leq 32p_{11}\eta^{2}$. Since $m\geq\frac{400\log(1/p^{2}\gamma)}{p^{2}\gamma}$, we have $m\geq\frac{384\log m}{p^{2}\gamma}\geq\frac{12\log m}{p_{11}\eta^{2}}$. Thus we get $m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\leq\exp(-\log m)$. When $m\geq\frac{400\log n}{p^{2}\phi(\gamma)}$, we get $n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\leq\exp(-\log n)$. Therefore, when $m\geq 400\left(\frac{\log(1/p^{2}\gamma)}{p^{2}\gamma}\vee\frac{\log n}{p^{2}\phi(\gamma)}\right)$, we have
$m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\leq\exp(-\log m),n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\leq\exp(-\log n).$
Therefore, $\sum_{k=1}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]\leq\frac{\exp(-\log m)}{1-\exp(-\log m)}+\frac{\exp(-\log n)}{1-\exp(-\log n)}$. We finish the proof.
4 Impossibility results
In this section, we present the impossibility results for the graph alignment problem. Under our proposed model, the alignment problem aims to recover the domain $S^{*}\subseteq V(G_{1})$, range $T^{*}\subseteq V(G_{2})$, and the mapping $\pi^{*}:S^{*}\mapsto T^{*}$. When equipped with the additional knowledge on $S^{*}$ and $T^{*}$, our problem can be reduced to recovery with full observations on smaller graphs, the reconstruction threshold for which is settled in [WXY22]. The lower bound therein remains valid when the number of correlated nodes is substituted with $m$. However, such reduction only proves tight in a limited number of regimes (see Proposition 4). We will establish the impossibility results for the remaining regimes by Fano’s method. Two main ingredients of Fano’s method are outlined as follows:
  • Construct a packing set ${\mathcal{M}}$ of the parameter space $\mathcal{S}_{n,m}$ such that any two distinct elements from ${\mathcal{M}}$ differ by a prescribed threshold. Specifically, in partial recovery, the overlap of each pair is less than $\delta$, which is equivalent to $\min_{\pi\neq\pi^{\prime}\in{\mathcal{M}}}d(\pi,\pi^{\prime})\geq(1-\delta)m$, while in exact recovery ${\mathcal{M}}=\mathcal{S}_{n,m}$. The cardinality of ${\mathcal{M}}$ measures the complexity of the parameter space under the target metric.
  • Choose the uniform prior on $\pi^{*}$ over ${\mathcal{M}}$ and upper bound the mutual information $I(\pi^{*};G_{1},G_{2})$. Given $\pi^{*}$, the conditional distribution of the observed graphs $(G_{1},G_{2})$ is specified in Definition 3. For the mutual information, let $\mathcal{P}$ denote the joint distribution of $(G_{1},G_{2})$ and $\mathcal{Q}$ be any distribution over $(G_{1},G_{2})$, then
    $I(\pi^{*};G_{1},G_{2})=\mathbb{E}_{\pi^{*}}[D(\mathcal{P}_{G_{1},G_{2}|\pi^{*}}\|\mathcal{P}_{G_{1},G_{2}})]\leq\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}}).$ (11)
    The impossibility results follows if $I(\pi^{*};G_{1},G_{2})\leq c\log|{\mathcal{M}}|$ for some small constant $c$.
Let ${\mathcal{M}}_{\delta}$ denote a packing set under the overlap threshold $\delta$. The size of ${\mathcal{M}}_{\delta}$ follows from the standard volume argument [PW22, Theorem 27.3]. For $r\in[m]$, let $B(\pi,r)\triangleq\{\pi^{\prime}:d(\pi,\pi^{\prime})\leq r\}$ denote the ball of radius $r$ centered at $\pi$. Then we have
$|{\mathcal{M}}_{\delta}|\geq\frac{|\mathcal{S}_{n,m}|}{\max_{\pi}|B(\pi,(1-\delta)m-1)|}\geq\frac{|\mathcal{S}_{n,m}|}{\max_{\pi}|B(\pi,(1-\delta)m)|}.$
It remains to evaluate the cardinality of $\mathcal{S}_{n,m}$ and upper bound the volume of the ball under our distance metric $d$. It is straightforward to obtain that $|\mathcal{S}_{n,m}|=\binom{n}{m}^{2}m!.$ Let $k=\delta m$. Note that all elements from $B(\pi,m-k)$ have at least $k$ common mappings. To upper bound $|B(\pi,m-k)|$, we first choose $k$ elements from the domain of $\pi$ and map to the same value as $\pi$, and the remaining domain and range of size $m-k$ and the mapping are selected arbitrarily. We get $|B(\pi,m-k)|\leq\tbinom{m}{k}\tbinom{n-k}{m-k}^{2}(m-k)!.$ Consequently,
$|{\mathcal{M}}_{\delta}|\geq\frac{\binom{n}{m}^{2}m!}{\binom{m}{k}\binom{n-k}{m-k}^{2}(m-k)!}=\left(\frac{\binom{n}{k}}{\binom{m}{k}}\right)^{2}k!>\left(\frac{n^{2}k}{e^{3}m^{2}}\right)^{k}\geq\left(\frac{\delta n}{e^{3}}\right)^{k},$ (12)
where we use the inequalities that $(\frac{n}{k})^{k}\leq\binom{n}{k}<(\frac{en}{k})^{k}$ and $k!\geq(k/e)^{k}$. Fano’s method provides a lower bound on the Bayesian risk when $\pi$ is uniformly distributed over ${\mathcal{M}}_{\delta}$, which further lower bound the minimax risk. The above argument also yields a lower bound when $\pi$ is uniform over $\mathcal{S}_{n,m}$ via generalized Fano’s inequality [BHM12, Lemma 20]. The following propositions provide lower bounds for $m$ for partial recovery and exact recovery, and thus prove the lower bounds in Theorems 1 and 2.
Proposition 3 (Lower bound for partial recovery).
For any $\delta\in(0,1)$, if $m\leq\frac{c\log n}{p^{2}\phi(\gamma)}$, then for any estimator $\hat{\pi}$,
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\geq 1-\frac{13c}{\delta}.$
Proof.
For any $\pi$ with domain $S$ and range $T$ such that $|S|=|T|=m$, arbitrarily pick a bijection $\sigma:V(G_{1})\mapsto V(G_{2})$ such that $\sigma|_{S}=\pi$. Then, the conditional distribution $\mathcal{P}_{G_{1},G_{2}|\pi}$ can be factorized into
$\mathcal{P}_{G_{1},G_{2}|\pi}=\prod_{e\in\tbinom{S}{2}}P(e,\pi(e))\prod_{e\in\tbinom{V(G_{1})}{2}\backslash\tbinom{S}{2}}Q(e,\sigma(e)),$
where $P\sim\mathrm{Bern}(p,p,\rho)$ and $Q\sim\mathrm{Bern}(p,p,0)$. Pick $\mathcal{Q}$ in (11) to be an auxiliary null model under which $G_{1}$ and $G_{2}$ are independent with the same marginal as $\mathcal{P}$. Then, $\mathcal{Q}_{G_{1},G_{2}}$ can be factorized into
$\mathcal{Q}_{G_{1},G_{2}}=\prod_{e\in\tbinom{S}{2}}Q(e,\pi(e))\prod_{e\in\tbinom{V(G_{1})}{2}\backslash\tbinom{S}{2}}Q(e,\sigma(e)).$
The KL-divergence between the product measures $\mathcal{P}_{G_{1},G_{2}|\pi}$ and $\mathcal{Q}_{G_{1},G_{2}}$ can be expressed as
$D\left(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}}\right)=\tbinom{m}{2}D(P\|Q)$
for any $\pi:S\mapsto T$ with $|S|=|T|=m$. Applying Lemma 5, we obtain
$\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq 25\tbinom{m}{2}p^{2}\phi(\gamma).$ (13)
Applying generalized Fano’s inequality [BHM12, Lemma 20] with (12) and (13), we obtain
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{25\tbinom{m}{2}p^{2}\phi(\gamma)}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{13c}{\delta},$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$. ∎
Proposition 4 (Lower bound for exact recovery).
For any $c\in(0,1)$ and any estimator $\hat{\pi}$, there exists constant $c_{3}$ only depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)\right)$,
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\geq 1-c,$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
Proof.
We first apply the reduction argument. With the additional information on the domain and range of $\pi^{*}$, our problem can be reduced to the reconstruction of mapping as in [WXY22]. Applying the lower bound in [WXY22, Theorem 4], for a fixed $\epsilon\in(0,1)$, when $m(\sqrt{p_{00}p_{11}}-\sqrt{p_{01}p_{10}})^{2}\leq(1-\epsilon)\log m$, we have $\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\geq 1-o(1)$ for any estimator $\hat{\pi}$. Note that $(\sqrt{p_{00}p_{11}}-\sqrt{p_{01}p_{10}})^{2}\asymp p^{2}(\gamma\wedge\gamma^{2})\asymp(\rho^{2})\wedge(\rho p).$ Therefore, when
$m\lesssim\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\log\left(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\right),$ (14)
we have $\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\geq 1-o(1)$. Applying Proposition 3 with $\delta=1/2$ yields that, when
$m\lesssim\frac{\log n}{p^{2}\phi(\gamma)},$ (15)
we have $\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\geq 1-c$ for $c\in(0,1)$.
When $\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\asymp n$, by (14), exact recovery is impossible, even when $m=n$. Next we consider the regime that $\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\lesssim n$. When $\gamma\leq 1$, we have $p^{2}(\gamma\wedge\gamma^{2})=p^{2}\gamma^{2}\asymp p^{2}\phi(\gamma)$, and thus
$\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\log\left(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\right)\lesssim\frac{\log n}{p^{2}\phi(\gamma)}.$
When $\gamma\geq 1$, $\gamma\wedge\gamma^{2}=\gamma$. By comparing (14) and (15), we derive that exact recovery is impossible if $m\lesssim\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)$. ∎
5 Discussion and future directions
This paper proposes the partially correlated Erdős-Rényi graphs model, wherein a pair of induced subgraphs with a certain size are correlated. We investigate the optimal information-theoretic threshold for recovering the latent correlated subgraphs and the hidden vertices correspondence under our new model. In comparison with prior work on correlated Erdős-Rényi graphs model, the additional challenge arises from the unknown location of the correlated subsets. For a candidate mapping $\pi$ whose domain may include both correlated and ambient subgraphs, we extend the classical notion of functional digraph to formally describe the correlation structure among the edges. We observe from the correlated functional digraph that the independent components consist of cycles and paths. The graphical representation may be of independent interest for general models.
There are many problems to be further investigated under our proposed model:
  • Refined results. The results in the paper could be further refined in various ways, such as deriving the sharp constants and characterizing the optimal scaling in terms of the fraction $\delta$ in partial recovery.
  • Efficient algorithms. It is of interest to investigate the polynomial-time algorithms and identify the computational hardness under our model. More efficient algorithms are also desirable when the signal is stronger.
  • Graph sampling. One motivation of the paper stems from graph sampling as discussed in Section 1.2. The sampled subgraphs are partially correlated, where the size of correlated subsets is a random variable depending on the sampling methods. Thus, it is natural to ask about the sample size needed for reliable recovery.
  • Correlation test. The correlation test problem under our model is also highly relevant. It is interesting to find out whether the detection problem is strictly easier than recovery, both in terms of the information thresholds and algorithmic developments.
Appendix A Proof of Lemmas
A.1 Proof of Lemma 1
We first evaluate the moment generating function for paths. Consider a path $P$ of size $\ell$ denoted by $\langle e_{1}e_{2}\dots e_{\ell}\rangle$ as illustrated in Figure 3. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq{\mathbf{1}_{\left\{{e_{i}\in E(G_{1})}\right\}}}$ and $B_{i}\triangleq{\mathbf{1}_{\left\{{\pi(e_{i})\in E(G_{2})}\right\}}}$. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$. By definition (5),
$|P\cap E({\mathcal{H}}_{\pi})|=\sum_{i=1}^{\ell}{\mathbf{1}_{\left\{{e_{i}\in E(G_{1})}\right\}}}{\mathbf{1}_{\left\{{\pi(e_{i})\in E(G_{2})}\right\}}}=\sum_{i=1}^{\ell}A_{i-1}B_{i}.$
[…]
Figure 3: Illustration of a path of size $\ell$.
For the sake of notational simplicity, we introduce an auxiliary random variable $B_{0}$ that is correlated with $A_{0}$ such that $(A_{0},B_{0})\sim\mathrm{Bern}(p,p,\rho)$. Then
$m_{\ell}\triangleq\mathbb{E}[e^{t|P\cap E({\mathcal{H}}_{\pi})|}]$ $=\mathbb{E}\left[\mathbb{E}\left[\prod_{i=1}^{\ell}e^{tA_{i-1}B_{i}}|B_{0}\dots B_{\ell}\right]\right]=\mathbb{E}\left[\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}B_{i}}|B_{i-1}B_{i}\right]\right]$
$=\sum_{b_{0},\dots,b_{\ell}\in\{0,1\}}\prod_{i=0}^{\ell}\mathbb{P}[B_{i}=b_{i}]\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right].$ (16)
Define $M(b_{i-1},b_{i})\triangleq\mathbb{P}[B_{i}=b_{i}]\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right]$ for $b_{i-1},b_{i}\in\{0,1\}$ and a matrix
$M\triangleq\begin{bmatrix}M(0,0)&M(0,1)\\ M(1,0)&M(1,1)\end{bmatrix}=\begin{bmatrix}\bar{p}&(\bar{p}+p_{01}(e^{t}-1))p/\bar{p}\\ \bar{p}&p+p_{11}(e^{t}-1)\end{bmatrix},$
where $\bar{p}=1-p$. Recall that $\mathbb{P}[B_{i}=1]=p$. Then we obtain that
$m_{\ell}=\sum_{b_{0},\dots,b_{\ell}\in\{0,1\}}\mathbb{P}[B_{0}=b_{0}]M(b_{0},b_{1})\dots M(b_{\ell-1},b_{\ell})=[\bar{p},p]M^{\ell}\begin{bmatrix}1\\ 1\end{bmatrix}.$
The trace and determinant of $M$ is
$T\triangleq\mathrm{Tr}(M)=1+p_{11}(e^{t}-1),\quad D\triangleq\det(M)=\rho p\bar{p}(e^{t}-1)>0.$
Since $D<p_{11}(e^{t}-1)$, the discriminant is $T^{2}-4D>0$. Hence, the matrix $M$ has two distinct eigenvalues denoted by $\lambda_{1}>\lambda_{2}>0$, and the general term of $m_{\ell}$ is
$m_{\ell}=\alpha\lambda_{1}^{\ell}+\beta\lambda_{2}^{\ell}.$ (17)
The coefficients $\alpha$ and $\beta$ can be determined via the first two terms $m_{0}=1$ and $m_{1}$. Then we get
$m_{\ell}=\left(\frac{1}{2}+\frac{2m_{1}-T}{2\sqrt{T^{2}-4D}}\right)\lambda_{1}^{\ell}+\left(\frac{1}{2}-\frac{2m_{1}-T}{2\sqrt{T^{2}-4D}}\right)\lambda_{2}^{\ell}.$
Furthermore, by plugging $m_{1}=1+p^{2}(e^{t}-1)$, we get $T-m_{1}=D$ and thus $m_{1}(T-m_{1})>D$, which is equivalent to $|2m_{1}-T|<\sqrt{T^{2}-4D}$. Therefore, both coefficients $\alpha,\beta\in(0,1)$.
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq{\mathbf{1}_{\left\{{e_{i}\in E(G_{1})}\right\}}}$ and $B_{i}\triangleq{\mathbf{1}_{\left\{{\pi(e_{i})\in E(G_{2})}\right\}}}$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$ for $i=0,\dots,\ell-1$.
[…]
Figure 4: Illustration of a cycle of size $\ell$.
Following a similar argument as (16), we have
$\tilde{m}_{\ell}\triangleq\mathbb{E}[e^{t|C\cap E({\mathcal{H}}_{\pi})|}]$ $=\sum_{b_{1},\dots,b_{\ell}=b_{0}\in\{0,1\}}\prod_{i=1}^{\ell}\mathbb{P}[B_{i}=b_{i}]\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right]$
$=\sum_{b_{1},\dots,b_{\ell}=b_{0}\in\{0,1\}}M(b_{0},b_{1})M(b_{1},b_{2})\dots M(b_{\ell-1},b_{0}).$
Applying the eigenvalue decomposition of $M$ again, we obtain that
$\tilde{m}_{\ell}=\mathrm{Tr}(M^{\ell})=\lambda_{1}^{\ell}+\lambda_{2}^{\ell}.$ (18)
By definition, $\kappa^{{\mathsf{P}}}_{\ell}(t)=\log m_{\ell}$ and $\kappa^{{\mathsf{C}}}_{\ell}(t)=\log\tilde{m}_{\ell}$. To upper bound the cumulants, it suffices to consider $m_{\ell}$ and $\tilde{m}_{\ell}$. In (17), we have $\alpha,\beta\in(0,1)$ and $\lambda_{1}>\lambda_{2}>0$. By monotonicity, it follows that $m_{\ell}\leq\tilde{m}_{\ell}$ and thus
$\kappa^{{\mathsf{P}}}_{\ell}(t)\leq\kappa^{{\mathsf{C}}}_{\ell}(t).$
For $x\in\mathbb{R}^{n}$ and $\ell\geq 2$, we have $\|{x}\|_{\ell}\leq\|{x}\|_{2}\leq\|{x}\|_{1}$. It follows from the formula of $\tilde{m}_{\ell}$ in (18) that $\tilde{m}_{\ell}^{1/\ell}\leq\tilde{m}_{2}^{1/2}\leq\tilde{m}_{1}$. Equivalently,
$\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\leq\kappa^{{\mathsf{C}}}_{1}(t),\quad\kappa^{{\mathsf{C}}}_{\ell}(t)\leq\frac{\ell}{2}\kappa^{{\mathsf{C}}}_{2}(t)\quad\forall~{}\ell\geq 2.$
The last inequality $2\kappa^{{\mathsf{P}}}_{1}(t)\leq\kappa^{{\mathsf{C}}}_{2}(t)$ follows by comparing the explicit formula $\kappa^{{\mathsf{P}}}_{1}(t)=\log(1+p^{2}(e^{t}-1))$ with $\kappa^{{\mathsf{C}}}_{2}(t)$ in (7) and using $p_{11}\geq p^{2}$.
Finally, since the summands over different connected components are independent, it follows that
$\log\mathbb{E}\left[e^{t|{\mathcal{E}}\cap E({\mathcal{H}}_{\pi})|}\right]$ $=\sum_{P\in{\mathcal{P}}}\kappa_{|P|}^{{\mathsf{P}}}(t)+\sum_{C\in{\mathcal{C}}}\kappa_{|C|}^{{\mathsf{C}}}(t)$
$\leq\sum_{P\in{\mathcal{P}}}\frac{|P|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+\sum_{C\in{\mathcal{C}}:|C|\geq 2}\frac{|C|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+\sum_{C\in{\mathcal{C}}:|C|=1}\kappa_{1}^{{\mathsf{C}}}(t)$
$=\frac{|{\mathcal{E}}|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+|\{C\in{\mathcal{C}}:|C|=1\}|\left(\kappa_{1}^{{\mathsf{C}}}(t)-\frac{1}{2}\kappa_{2}^{{\mathsf{C}}}(t)\right),$
where the last equality used fact that $|{\mathcal{E}}|=\sum_{P\in{\mathcal{P}}}|P|+\sum_{C\in{\mathcal{C}}}|C|$.
Remark 1.
We have two bounds for large $\ell$ in Lemma 1, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha\lambda_{1}^{\ell}+\beta\lambda_{2}^{\ell}\right)\leq\frac{1}{\ell}\log(\lambda_{1}^{\ell}+\lambda_{2}^{\ell})$, where $10<\beta<\alpha<1,\alpha+\beta=$ 1 and $\lambda_{1}>\lambda_{2}>0$. Consequently, $\lambda_{1}-\frac{\log 2}{\ell}\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq\lambda_{1}+\frac{\log 2}{\ell}.$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $\|x\|_{\ell}\leq\|x\|_{2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
A.2 Proof of Lemma 2
By Lemma 1,
$\log\mathbb{E}\left[e^{t|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|}\right]\leq\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$
where $L$ denotes the number of self-loops. The self-loop for $e$ only happens when $\pi(e)=\pi^{*}(e)$. For $uv\in\tbinom{V(G_{1})}{2}\backslash\tbinom{F_{\pi}}{2}$, by the definition of $F_{\pi}$, $\pi(u)\neq\pi^{*}(u)$ or $\pi(v)\neq\pi^{*}(v)$. Therefore, $\pi(uv)=\pi^{*}(uv)$ implies that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$. Since $d(\pi^{*},\pi)=k$, we must have $L\leq\frac{k}{2}$.
Applying the formulas (6) and (7) and the fact that $p_{11}\leq p$, we obtain
$\kappa^{{\mathsf{C}}}_{2}(t)$ $\leq\log\left(1+2p^{2}(e^{t}-1)+p^{2}(e^{t}-1)^{2}\right)$
$=\log\left(1+p^{2}(e^{2t}-1)\right)\leq p^{2}(e^{2t}-1)$
and
$\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)$ $=\frac{1}{2}\log\left[1+\frac{2(p_{11}-p^{2})}{p_{11}^{2}(e^{t}-1)+2p^{2}+(e^{t}-1)^{-1}}\right]$
$\overset{\mathrm{(a)}}{\leq}\frac{1}{2}\log\left[1+\frac{2(p_{11}-p^{2})}{2(p_{11}+p^{2})}\right]\overset{\mathrm{(b)}}{\leq}\frac{\gamma}{2(\gamma+2)},$
where $\mathrm{(a)}$ is because $x+x^{-1}\geq 2$ for any $x>0$ and $\mathrm{(b)}$ is because $\log(1+x)\leq x$ for any $x\geq 0$. Therefore, we get
$\log\mathbb{E}\left[e^{t|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|}\right]\leq\frac{|{\mathcal{E}}_{\pi}|}{2}p^{2}(e^{2t}-1)+\frac{k\gamma}{4(\gamma+2)}.$
For any $t>0$, by the Chernoff bound,
$\mathbb{P}[|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|\geq\tau_{k}]\leq\exp\left(-t\tau_{k}+\frac{|{\mathcal{E}}_{\pi}|}{2}p^{2}(e^{2t}-1)+\frac{k\gamma}{4(2+\gamma)}\right).$
Pick $t=\frac{1}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)$. Then $t>0$ by the assumption $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$. We obtain
$\mathbb{P}\left[|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|\geq\tau_{k}\right]\leq\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}+\frac{k\gamma}{4(2+\gamma)}\right).$
A.3 Proof of Lemma 3
Lemma 3.
Recall that $\phi(\gamma)=(1+\gamma)\log(1+\gamma)-\gamma$ and $\eta,\gamma>0$. If $\eta\leq\frac{\gamma}{4(1+\gamma)}$, then $(1+\gamma)(1-\eta)>1$ and
$\phi\left[(1-\eta)(1+\gamma)-1\right]\geq\frac{1}{4}\phi(\gamma).$ (19)
Proof.
We note that $(1+\gamma)(1-\eta)\geq 1+\gamma-\frac{\gamma}{4}>1$ and
$\phi\left[(1-\eta)(1+\gamma)-1\right]=$ $(1+\gamma)(1-\eta)\log\left[(1+\gamma)(1-\eta)\right]-\left[(1+\gamma)(1-\eta)-1\right]$
$=$ $(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+(1+\gamma)(1-\eta)\log(1-\eta)+\eta$
$\geq$ $(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+(1+\gamma)(-\eta)+\eta$
$=$ $(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]-\eta\gamma,$
where the last inequality is due to the fact that $(1-x)\log(1-x)+x\geq 0$ for any $10<x<$ 1 and $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. Since $\log(1+\gamma)-\frac{\gamma}{1+\gamma}\geq\frac{\gamma^{2}}{2(1+\gamma)^{2}}$, we have $\eta\gamma\leq\frac{\gamma^{2}}{4(1+\gamma)}\leq\frac{1}{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$. Therefore,
$\phi\left[(1-\eta)(1+\gamma)-1\right]\geq$ $(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]-\eta\gamma$
$\geq$ $\left(\frac{1}{2}-\eta\right)\phi(\gamma)\geq\frac{1}{4}\phi(\gamma),$
where the last inequality is due to $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. ∎
A.4 Proof of Lemma 4
Lemma 4.
For binary entropy function $h(x)=-x\log x-(1-x)\log(1-x)$, $\phi(x)=(1+x)\log(1+x)-x$ and any constant $\delta\in(0,1)$, when $m\geq\log n$,
$\sum_{k=\delta m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]\leq\left(\log n\right)^{-1+o(1)}$
Proof.
We note that
$\sum_{k=\delta m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]$ $\leq\sum_{k=1}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]$
$\overset{\mathrm{(a)}}{\leq}2\sum_{1\leq k\leq\frac{m}{2}}\exp\left[-k\log\left(\frac{m}{k}\right)\right]$
$\leq 2\sum_{1\leq k\leq 2\log m}\exp\left[-k\log\left(\frac{m}{k}\right)\right]+2\sum_{2\log m<k\leq\frac{m}{2}}2^{-k}$
$\leq 2\cdot\exp\left(-\log m\right)\cdot(2\log m)+2\cdot 2^{-2\log m}\overset{\mathrm{(b)}}{\leq}(\log n)^{-1+o(1)},$
where $\mathrm{(a)}$ is because $h(x)=h(1-x)$ and $h(x)\geq-x\log x$ and $\mathrm{(b)}$ is because $m\geq\log n$. ∎
A.5 Proof of Lemma 5
Lemma 5.
For $P\sim\mathrm{Bern}(p,p,\rho)$ and $Q\sim\mathrm{Bern}(p,p,0)$, the KL-divergence between $P$ and $Q$ can be upper bounded by:
$D(P\|Q)\leq 25p^{2}\phi(\gamma).$
Proof.
By direct calculation,
$D(P\|Q)=\sum_{\{a,b\}\in\{0,1\}}p_{ab}\log\left[\frac{p_{ab}}{p^{a+b}(1-p)^{2-a-b}}\right]$
$=$ $\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]+2p(1-p)(1-\rho)\log(1-\rho)$
$+$ $\left[(1-p)^{2}+\rho p(1-p)\right]\log\left(1+\frac{\rho p}{1-p}\right)$
$\leq$ $\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]+2p(1-p)(1-\rho)\cdot(-\rho)$
$+$ $\left[(1-p)^{2}+\rho p(1-p)\right]\cdot\frac{\rho p}{1-p}$
$=$ $\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]-\rho p(1-p)+\rho^{2}\left[2p(1-p)+p^{2}\right].$
Since $\log(1+x)\geq\frac{x}{x+1}+\frac{x^{2}}{2(x+1)^{2}}$ for any $x\geq 0$, we get $p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\geq\frac{p^{2}\gamma^{2}}{2(\gamma+1)}$. When $\gamma<3$, since $\frac{p^{2}\gamma^{2}}{2(\gamma+1)}\geq\frac{p^{2}\gamma^{2}}{8}=\frac{\rho^{2}(1-p)^{2}}{8}\geq\frac{\rho^{2}\left[2p(1-p)+p^{2}\right]}{24}$ for $0<p\leq\frac{1}{2}$, we get
$D(P\|Q)\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right]\leq 25p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right].$
When $\gamma\geq 3$, since $p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\geq p^{2}\gamma(\log 4-1)=(\log 4-1)\rho p(1-p)$ and $\rho^{2}\left[2p(1-p)+p^{2}\right]\leq 3\rho p(1-p)$, we get
$D(P\|Q)$ $\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right]$
$\leq\left(\frac{3}{\log 4-1}+1\right)p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\leq 25p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right].$
Therefore, we get $D(P\|Q)\leq 25p^{2}\phi(\gamma)$. ∎
A.6 Proof of Lemma 6
Lemma 6 (Chernoff’s inequality for Binomials).
Suppose $\xi\sim\mathrm{Bin}(n,p)$, denote $\mu=np$, then
$\mathbb{P}\left[\xi\geq(1+\delta)\mu\right]\leq\exp\left\{-\mu\left[(1+\delta)\log(1+\delta)-\delta\right]\right\},$ (20)
$\mathbb{P}\left[\xi\leq(1-\delta)\mu\right]\leq\exp\left(-\frac{\delta^{2}\mu}{2}\right).$ (21)
We also have
$\mathbb{P}\left[\xi\geq(1+\delta)\mu\right]\leq\exp\left(-\frac{\delta\mu}{2+\delta}\right).$ (22)
Proof.
By Theorems 4.4 and 4.5 in [MU05] we have (20) and (21). Since $(1+\delta)\log(1+\delta)-\delta\geq\frac{\delta^{2}}{2+\delta}$, we obtain (22) from (20). ∎