### v2

# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

Abstract
This paper studies the problem of recovering hidden vertex correspondences between two correlated random graphs. We introduce the partially correlated Erdős-Rényi model and the partially correlated Gaussian Wigner model, where a pair of induced subgraphs is correlated. We investigate the information-theoretic thresholds for recovering these latent correlated subgraphs and their hidden vertex correspondences. For the partially correlated Erdős-Rényi model, we establish the optimal rate for partial recovery: above this threshold, a positive fraction of vertices can be correctly matched, while below it, matching any positive fraction is impossible. We also determine the optimal rate for exact recovery. In the partially correlated Gaussian Wigner model, the optimal rates for partial and exact recovery coincide. To prove the achievability results, we introduce correlated functional digraphs to partition the edges and bound error probabilities using lower-order cumulant generating functions. Our impossibility results rely on a generalized Fano’s inequality and the recovery thresholds for correlated Erdős-Rényi graphs.
Keywords— Graph alignments, information-theoretic thresholds, Erdős-Rényi model, Gaussian Wigner model, partial recovery, exact recovery
1 Introduction
Recently, there has been a surge of interest in the problems of detecting graph correlations and recovering the alignments of two correlated graphs. These questions have emerged across various domains. For instance, in social networks, determining the similarity between friendship networks across different platforms has garnered attention [NS08, NS09]. In the realm of computer vision, the identification of whether two graphs represent the same object holds significant importance in pattern recognition and image processing [BBM05, CSS06]. In computational biology, the representation of biological networks as graphs aids in understanding and quantifying their correlation [SXB08, VCL+15]. Furthermore, in natural language processing, the task of determining whether a given sentence can be inferred from the text directly relates to graph matching problems [HNM05].
Numerous graph models exist, with the Erdős-Rényi random graph model being a prominent example, as proposed by [ER59] and [Gil59]:
Definition 1 (Erdős-Rényi graph).
The Erdős-Rényi random graph is the graph on $n$ vertices where each edge connects with probability $10<p<$ 1 independently. Let $\mathcal{G}(n,p)$ denote the distribution of Erdős-Rényi random graphs with $n$ vertices and edge connecting probability $p$.
While there are inherent disparities between the Erdős-Rényi graph and networks derived from real-world scenarios, comprehensively understanding the Erdős-Rényi graphs remains profoundly significant. This understanding serves as a pivotal step in transitioning from solving detection and matching problems on Erdős-Rényi graphs to addressing challenges inherent in practical applications. The graph alignment problem entails identifying latent vertex correspondences between two graphs based on their structures. Following [PG11], for two random graphs $G_{1},G_{2}$, a common graph model is the correlated Erdős-Rényi model. For a weighted graph $G$ with vertex set $V(G)$ and edge set $E(G)$, the weight associated with each edge $uv$ is denoted as $\beta_{uv}(G)$ for any $u,v\in V(G)$. For an unweighted graph $G$, we define $\beta_{uv}(G)={\mathbf{1}_{\left\{{uv\in E(G)}\right\}}}$.
Definition 2 (Correlated Erdős-Rényi graphs).
Let $\pi$ denote a latent bijective mapping from $V(G_{1})$ to $V(G_{2})$. We say a pair of graphs $(G_{1},G_{2})$ is correlated Erdős-Rényi graphs if both marginal distributions are $\mathcal{G}(n,p)$ and each pair of edges $(\beta_{uv}(G_{1}),\beta_{\pi(u)\pi(v)}(G_{2}))$ for $u,v\in V(G_{1})$ follows the correlated bivariate Bernoulli distribution with correlation coefficient $\rho$.
We note that the edges in the Erdős-Rényi model are binary 0-1 random variables, where $\beta_{uv}(G_{1}),\beta_{\pi(u)\pi(v)}(G_{2})\in\left\{0,1\right\}$ for any $u,v\in V(G_{1})$. The Bernoulli-based Erdős-Rényi model, while useful for modeling binary relationships, can be limited when we aim to represent weighted edges or capture more complex dependencies between nodes. Another important model is the correlated Gaussian Wigner model proposed in [DMWX21] as a prototypical model for random graphs, where the edges follow Gaussian distributions.
Definition 3 (Correlated Gaussian Wigner model).
Let $\pi$ denote a latent bijective mapping from $V(G_{1})$ to $V(G_{2})$. We say a pair of graphs $(G_{1},G_{2})$ follows correlated Gaussian Wigner model if each pair of weighted edges $(\beta_{uv}(G_{1}),\beta_{\pi(u)\pi(v)}(G_{2}))$ follows bivariate normal distribution $\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for any vertices $u,v\in V(G_{1})$.
Given observations on $G_{1}$ and $G_{2}$ under the correlated Erdős-Rényi graphs model or correlated Gaussian Wigner model, the goal is to recover the latent vertex mapping $\pi$. Let $\mathrm{Supp}(\pi)$ denote the support set of $\pi$. To quantify the performance of an estimator $\hat{\pi}$, we consider the following two recovery criteria:
  • Partial recovery: given a constant $\delta\in(0,1)$, we say $\hat{\pi}$ succeeds for partial recovery if
    |$|\{v\in\mathrm{Supp}(\pi):\pi(v)=\hat{\pi}(v)\}|\geq\delta|\mathrm{Supp}(\pi)|.$ (1)
  • Exact recovery: we say $\hat{\pi}$ succeeds for exact recovery if
    $\pi(v)=\hat{\pi}(v),\quad\forall~{}v\in\mathrm{Supp}(\pi).$ (2)
The information-theoretic thresholds for partial and exact recoveries of $\pi$ under correlated Erdős-Rényi model and correlated Gaussian Wigner model have been extensively studied in the recent literature.
  • Erdős-Rényi model, partial recovery. In the sparse regime where $np$ and $\rho$ are constant, partial recovery is impossible [GML21]. It is shown in [HM23] that $np(p\vee\rho)\gtrsim\log\left(1+\frac{\rho}{p}\right)\vee 1$ suffices for partial recovery, while $n\gtrsim d(p+\rho-p\rho\|p)\log n$ is necessary, where $d(p\|q)$ denotes the Kullback–Leibler (KL) divergence between Bernoulli distributions with mean $p$ and $q$, respectively. The recent work [WXY22] settled the sharp threshold for dense graphs with $\frac{p}{p\vee\rho}=n^{-o(1)}$ and the thresholds within a constant factor for sparse ones with $\frac{p}{p\vee\rho}=n^{-\Omega(1)}$. For the sparse case, a sharp threshold has been proven when $\frac{p}{p\vee\rho}=n^{-\alpha+o(1)}$ for $\alpha\in(0,1]$ in [DD23b].
  • Erdős-Rényi model, exact recovery. Based on the properties of the intersection graph under a permutation $\pi$, it is shown in [CK16, CK17] that the Maximal Likelihood Estimator (MLE) achieves exact recovery and establishes an information-theoretical lower bound with a gap of $\omega(1)$. The results are sharpened by [WXY22] where the sharp threshold for exact recovery are derived.
  • Gaussian Wigner model. It is shown in [Gan22] that if $n\rho^{2}\geq(4+\epsilon)\log n$ for any constant $\epsilon>0$, then the MLE achieves exact recovery; if instead $n\rho^{2}\leq(4-\epsilon)\log n$, then exact recovery is impossible. The results are strengthened by [WXY22] by showing that even partial recovery is impossible under the same condition.
While numerous studies have extensively investigated recovery procedures in correlated Erdős-Rényi and correlated Gaussian Wigner models, it is however imperative to recognize that, in real-world applications, many nodes in one graph may not have corresponding counterparts in the other graph, leading to incomplete or misaligned structural information. To offer a resolution to this concern, we propose the following models where only a subset of the nodes between the two graphs are correlated.
Definition 4 (Partially correlated Erdős-Rényi graphs).
Let $S^{*}\subseteq V(G_{1})$ be a latent subset of vertices and $\pi^{*}:S^{*}\mapsto V(G_{2})$ be a latent injective mapping. We say a pair of graphs $(G_{1},G_{2})$ is partially correlated Erdős-Rényi graphs if both marginal distributions are $\mathcal{G}(n,p)$ and each pair of weighted edges $(\beta_{uv}(G_{1}),\beta_{\pi^{*}(u)\pi^{*}(v)}(G_{2}))$ for $u,v\in S^{*}$ follows the correlated bivariate Bernoulli distribution with correlation coefficient $\rho$.
Definition 5 (Partially correlated Gaussian Wigner model).
Let $S^{*}\subseteq V(G_{1})$ be a latent subset of vertices and $\pi^{*}:S^{*}\mapsto V(G_{2})$ be a latent injective mapping. We say a pair of graphs $(G_{1},G_{2})$ follows partially correlated Gaussian Wigner model if the marginal distribution of each edge in both graphs is standard normal, and for $u,v\in S^{*}$, the pair $(\beta_{uv}(G_{1}),\beta_{\pi^{*}(u)\pi^{*}(v)}(G_{2}))$ follows bivariate normal distribution with correlation coefficient $\rho$.
Let $G[S]$ denote the induced subgraph of $G$ with vertex set $S\subseteq V(G_{1})$. For the partially models in Definitions 4 and 5, given $S^{*}\subseteq V(G_{1})$ and the range of $\pi^{*}$ denoted by $T^{*}=\pi^{*}(S^{*})\subseteq V(G_{2})$, the induced subgraphs $G_{1}[S^{*}]$ and $G_{2}[T^{*}]$ follow correlated Erdős-Rényi model and correlated Gaussian Wigner model on $m$ vertices, respectively. Specifically, the case $S^{*}=V(G_{1})$ reduces to correlated Erdős-Rényi model and correlated Gaussian Wigner model in Definitions 2 and 3.
In this paper, we investigate the information-theoretic thresholds for recovering the set of correlated nodes $S^{*}$ and the mapping $\pi^{*}$. For notational simplicity, we also refer to the problem as recovering $\pi^{*}$ while keeping $S^{*}$ implicit as the domain of $\pi^{*}$. The success criteria are given by (1) and (2), where $\mathrm{Supp}(\pi)=S^{*}$. Owing to the potential inconsistency between the domain of $\pi^{*}$ and that of the estimator $\hat{\pi}:\hat{S}\mapsto V(G_{2})$, we define their overlap by
$\mathrm{overlap}(\pi^{*},\hat{\pi})\triangleq\frac{|v\in S^{*}\cap\hat{S}:\pi^{*}(v)=\hat{\pi}(v)|}{|S^{*}|}.$ (3)
With the notion of overlap, the success criteria are equivalent to
  • Partial recovery: $\hat{\pi}$ succeeds if $\mathrm{overlap}(\pi^{*},\hat{\pi})\geq\delta$ for a given constant $\delta\in(0,1)$;
  • Exact recovery: $\hat{\pi}$ succeeds if $\mathrm{overlap}(\pi^{*},\hat{\pi})=1$.
1.1 Main Results
In this subsection, we present the main results of the paper. We first introduce some notations for the presentation of the main theorems. Throughout the paper, we assume that $10<\rho<$ 1, $0<p\leq\frac{1}{2}$, and the cardinality $|S^{*}|=m$ is known. We denote the bivariate distribution of a pair of Bernoulli random variables with means $p_{1},p_{2}$, and correlation coefficient $\rho$ as $\mathrm{Bern}(p_{1},p_{2},\rho)$. Specifically, for $\mathrm{Bern}(p,p,\rho)$, we denote the following probability mass function:
$p_{11}\triangleq p^{2}+\rho p(1-p),\quad p_{10}=p_{01}\triangleq(1-\rho)p(1-p),\quad p_{00}\triangleq(1-p)^{2}+\rho p(1-p).$ (4)
  • In the Erdős-Rényi model, a pair of correlated edges
    $\left(\beta_{uv}(G_{1}),\beta_{\pi^{*}(u)\pi^{*}(v)}(G_{2})\right)\sim\mathrm{Bern}(p,p,\rho).$
    Specifically, two correlated edges are both present with probability $p_{11}$, whereas two independent edges are both present with probability $p^{2}$. The relative signal strength is quantified by $\gamma\triangleq\frac{p_{11}}{p^{2}}-1=\frac{\rho(1-p)}{p}$. This reparametrization of the correlation coefficient is crucial in determining the fundamental limits of the graph alignment problem.
  • In the Gaussian Wigner model, a pair of correlated edges consists of two standard Gaussian random variables with correlation coefficient $\rho\in(0,1)$:
    $(\beta_{uv}(G_{1}),\beta_{\pi^{*}(u)\pi^{*}(v)}(G_{2}))\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right).$
    Here, the relative signal strength is directly characterized by the correlation coefficient $\rho$.
In the Erdős-Rényi model, we assume $p\geq\frac{1}{n}$, as partial recovery is otherwise impossible [WXY22]. Define
$\phi(\gamma)\triangleq(1+\gamma)\log(1+\gamma)-\gamma,\quad\gamma\triangleq\frac{\rho(1-p)}{p},$ (5)
and let $\mathcal{S}_{n,m}$ denote the set of injective mappings $\pi:S\subseteq V(G_{1})\mapsto V(G_{2})$ with $|S|=m$. Our goal is to determine the minimum number of correlated nodes $m$ required for successful recovery of $\pi^{*}$. Next, we introduce our main theorems.
Theorem 1 (Erdős-Rényi model, partial recovery).
There exists an estimator $\hat{\pi}$ such that, for any constant $\delta\in(0,1)$ and $\pi^{*}\in\mathcal{S}_{n,m}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})\geq\delta\right]=1-o(1),$
when $m\geq\frac{c_{1}(\delta)\log n}{p^{2}\phi(\gamma)}$, where $c_{1}(\delta)$ is a constant depending on $\delta$.
Conversely, for any constant $c,\delta\in(0,1)$, there exists a constant $c_{2}(c,\delta)$ such that, when $m\leq\frac{c_{2}(c,\delta)\log n}{p^{2}\phi(\gamma)}$, for any estimator $\hat{\pi}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-c,$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
The possibility result is established in the minimax sense, while the impossibility result is under a Bayesian framework. Hence, the threshold applies to both minimax and Bayesian risks. Theorem 1 implies that for partial recovery, the threshold for the number of correlated nodes $m$ is of the order $\frac{\log n}{p^{2}\phi(\gamma)}$, beyond which partial recovery is possible and below which partial recovery is impossible. The dependency on the ambient graph order $n$ is logarithmic, while the scaling with respect to $p$ and $\rho$ is characterized by $\frac{1}{p^{2}\phi(\gamma)}$.
Theorem 2 (Erdős-Rényi model, exact recovery).
There exists an estimator $\hat{\pi}$ such that, for any $\pi^{*}\in\mathcal{S}_{n,m}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})=1\right]=1-o(1),$
when $m\geq C_{1}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\right)$, where $C_{1}$ is a universal constant.
Conversely, for any $c\in(0,1)$, there exists a constant $c_{3}$ only depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\right)$, for any estimator $\hat{\pi}$
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<1\right]\geq 1-c,$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
For exact recovery, Theorem 2 establishes the threshold for $m$ of the order $\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$. Under the weak signal regime where $\gamma=O(1)$, this rate coincides with that for partial recovery in Theorem 1. While the logarithmic scaling in $n$ is common to many other problems on random graphs, under the strong signal regime where $\gamma=\omega(1)$, Theorem 2 reveals a transition from $\frac{\log n}{p^{2}\phi(\gamma)}$ to $\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$ when $\log^{2}\frac{1}{p}-\log^{2}\frac{1}{\rho}\gtrsim\log n$. In this regime, the challenge is essentially the oracle recovery of mapping given the sets of correlated nodes $(S^{*},T^{*})$. See more discussions in Section 4.
Theorem 3 (Gaussian Wigner model).
For any $\rho\in(0,1)$, there exists an estimator $\hat{\pi}$ such that, for any $\pi^{*}\in\mathcal{S}_{n,m}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})=1\right]=1-o(1),$
when $m\geq C_{2}\left(\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1\right)$, where $C_{2}$ is a universal constant.
Conversely, for any constant $c,\delta\in(0,1)$, there exists a constant $c_{4}(c,\delta)$ such that, when $m\leq c_{4}(c,\delta)\left(\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1\right)$, for any estimator $\hat{\pi}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-c,$
where $\pi^{*}$ is uniformly distributed over $\mathcal{S}_{n,m}$.
Theorem 3 implies that in the Gaussian Wigner model, the thresholds for the number of correlated nodes $m$ are of the order $\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1$, beyond which exact recovery is possible and below which partial recovery is impossible. There is no gap between partial and exact recovery. Indeed, the estimation of $\pi^{*}$ can be divided into two steps: identifying the correlated nodes $(S^{*},T^{*})$ and recovering the mapping given $(S^{*},T^{*})$. Since the relative signal strength is characterized by the bounded correlation coefficient $\rho$, identifying the correlated nodes is always the major challenge, with no gap existing under the two recovery regimes.
Remark 1.
The optimal rates for partial and exact recoveries under both correlated Erdős-Rényi model and correlated Gaussian Wigner model are derived in [WXY22]. Our results applied with $S^{*}=V(G_{1})$ match the thresholds established in their work up to a constant factor. Furthermore, the lower bound $\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$ for exact recovery is derived from addressing the alignment problem for the subgraphs with the additional information on the domain and range of $\pi^{*}$, which applies the result in their work.
1.2 Related Work
Graph sampling.
When analyzing the properties of graphs, several challenges arise, such as limited data, high acquisition costs [SWM05], and incomplete knowledge of hidden structures [LF09, YML13, FH16]. Due to these challenges, graph sampling is a powerful approach for exploring graph structure. When data is sampled from two large networks, it is often unrealistic to assume full correlation among nodes in the sampled subgraphs. This naturally leads to partially correlated graphs. While the exact number of correlated nodes may be unknown, we usually have some estimate of the size. As a simplification, our model considers the case where the size is fixed. For the case where the size of correlated nodes is unknown, we leave it as our future work. See more details in Remark 4.
Efficient algorithms and computational hardness.
Numerous algorithms have been developed for the recovery problem. For example, percolation graph matching algorithm [YG13], subgraph matching algorithm [BCL+19], and local neighborhoods based algorithm [MX20]. However, these algorithms may be computationally inefficient. There are several polynomial-time algorithms for recovery, catering to different regimes correlation coefficients $\rho$. These include works by [BES80, Bol82, DCKG19, GM20, DMWX21, MRT23, MWXY23, DL23, MS24]. For instance, a polynomial-time algorithm for recovery is proposed in [MWXY23] by counting chandeliers when the correlation coefficient $\rho>\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant introduced in [Ott48]. Additionally, there exists an efficient iterative polynomial-time algorithm for sparse Erdős-Rényi graphs when the correlation coefficient is a constant [DL23].
It is postulated in [HS17, Hop18, KWB19] that the framework of low-degree polynomial algorithms effectively demonstrates computation hardness of detecting and recovering latent structures, and it bears similarities to sum-of-square methods [HKP+17, Hop18]. Based on the conjecture on the hardness of low-degree polynomial algorithms, it is shown in [MWXY24] that there is no polynomial-time test or matching algorithm when the correlation coefficient satisfies $\rho^{2}\leq\frac{1}{\text{polylog}(n)}$. Furthermore, the recent work [DDL23] showed computation hardness for detection and exact recovery when $p=n^{-1+o(1)}$ and the correlation coefficient $\rho<\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant, suggesting that several polynomial algorithms may be essentially optimal.
The maximal overlap is a test statistic which aims to identify the mapping that maximizes the edge correlation between two graphs. It is known that finding the maximal overlap is an instance of quadratic assignment problem (QAP) [PRW94], which is NP-hard to solve or to approximate [MMS10]. There are many studies aiming to detect or recover latent structures based on the maximal overlap statistics [CK16, CK17, BCL+19, MX20, DMWX21, WXY22, WXY23, HM23]. Finally, we mention that the recent work [DDG24] approximated the maximal overlap within a constant factor in polynomial-time for sparse Erdős-Rényi graphs, and [DGH23] established a sharp transition on approximating problem on the performance of online algorithms for dense Erdős-Rényi graphs. In this paper, we do not address polynomial algorithms or the computational hardness on partial models and leave these topics for future work.
Correlation detection
Besides the recent literature on the graph alignment problem, the correlation detection is another related topic. Given a pair of graphs, their correlation detection is formulated as a hypothesis testing problem, wherein the null hypothesis assumes independent random graphs, while the alternative assumes edge correlation under a latent permutation. A hypothesis testing model for correlated Erdős-Rényi graphs is proposed in [BCL+19]. The sharp threshold for dense Erdős-Rényi graphs and the threshold within a constant factor for sparse Erdős-Rényi graphs on this hypothesis testing model are established in [WXY23]. It is shown in [DD23a] that the sharp threshold for sparse Erdős-Rényi graphs can be derived by analyzing the densest subgraph. Additionally, a polynomial time algorithm for detection is also possible by counting trees when the correlation coefficient exceeds a constant value [MWXY24]. It is natural to ask whether the correlation can be detected when only a subsample from the graphs is collected. The probabilistic model is similar to the one present in the current paper, and we leave the exploration as our future work.
Other graph models.
Many properties of the correlated Erdős-Rényi model have been extensively investigated. However, the strong symmetry and tree-like structure inherent in this model distinguish it significantly from graph models encountered in practical applications. Therefore, it is crucial to explore more general graph models. One such model is inhomogeneous random graph model, where the edge connecting probability varies among edges in the graph [RS23, SPT23, DFW23]. Besides, geometric random graph model [WWXY22, SNL+23, BB24, GL24], planted cycle model [MWZ23, MWZ24], planted subhypergraph model [DMW23] and corrupt model [AH24] have also been subjects of recent studies.
1.3 Notations
For any $n\in\mathbb{N}$, let $[n]\triangleq\{1,2,\cdots,n\}$. For any $a,b\in\mathbb{R}$, let $a\wedge b=\min\{a,b\}$ and $a\vee b=\max(a,b)$. We use standard asymptotic notation: for two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$, if $a_{n}\leq Cb_{n}$ for some absolute constant $C$ and all $n$; $a_{n}=\Omega(b_{n})$ or $a_{n}\gtrsim b_{n}$, if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$ or $a_{n}\asymp b_{n}$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$.
For a given weighted graph $G$, let $V(G)$ denote its vertex set and $E(G)$ its edge set. For a set $V$, let $\binom{V}{2}\triangleq\{\{x,y\}:x,y\in V,x\neq y\}$ denote the collection of all subsets of $V$ with cardinality two. We write $uv$ to represent an edge $\{u,v\}$, and $\beta_{e}(G)$ for the weight of the edge $e$. For unweighted graphs $G$, we define $\beta_{uv}(G)={\mathbf{1}_{\left\{{uv\in E(G)}\right\}}}$. Let $\mathsf{{v}}(G)=|V(G)|$ denote the number of vertices in $G$, and $\mathsf{{e}}(G)=\sum_{e\in E(G)}\beta_{e}(G)$ the total weight of its edges. The induced subgraph of $G$ over a vertex set $V$ is denoted by $G[V]$. Given an injective mapping of vertices $\pi:S\subseteq V(G_{1})\mapsto V(G_{2})$, the induced injective mapping of edges is defined as $\pi^{{\mathsf{E}}}:\binom{S}{2}\mapsto\binom{V(G_{2})}{2}$, where $\pi^{{\mathsf{E}}}(uv)=\pi(u)\pi(v)$ for $u,v\in S$. For simplicity, we write $\pi(e)$ to denote $\pi^{{\mathsf{E}}}(e)$ when it is clear from the context.
2 Correlated functional digraph
A mapping from a set to itself can be graphically represented as a functional digraph (see, e.g., [Wes21, Definition 1.3.3]). Here, we extend this notion to mappings between distinct domain and range sets, where elements from the two sets are correlated. Although our focus in this section is on the mappings between the edges of $G_{1}$ and $G_{2}$, the graphical representation can be easily generalized to mappings between any two finite sets, such as sets of vertices.
Given a domain subset $S\subseteq V(G_{1})$, an injective function $\pi:S\mapsto V(G_{2})$, and a bivariate function $f:\mathbb{R}\times\mathbb{R}\mapsto\mathbb{R}$, we define the $f$-intersected graph ${\mathcal{H}}_{\pi}^{f}$ as
$V({\mathcal{H}}_{\pi}^{f})=V(G_{1}),\quad\beta_{e}(\mathcal{H}_{\pi}^{f})=\begin{cases}f\left(\beta_{e}(G_{1}),\beta_{\pi^{\mathsf{E}}(e)}(G_{2})\right),&\text{ if }e\in E(G_{1})\cap\binom{S}{2},\\ 0,&\text{ otherwise.}\end{cases}$
The weight in ${\mathcal{H}}_{\pi}^{f}$ represents the similarity score under $\pi$. To establish the possibility results, our estimator maximizes the total similarity score:
$\hat{\pi}(f)$ $\in\mathop{\text{argmax}}\limits_{\pi\in\mathcal{S}_{n,m}}\mathsf{{e}}(\mathcal{H}_{\pi}^{f})=\mathop{\text{argmax}}\limits_{\pi\in\mathcal{S}_{n,m}}\sum_{e\in E(G_{1})}\beta_{e}(\mathcal{H}_{\pi}^{f}).$ (6)
Specifically, we use $\hat{\pi}(f)$ with $f(x,y)=xy$ to maximize overlap, and $f(x,y)=-\frac{1}{2}(x-y)^{2}$ to minimize the mean-squared distance. For notational simplicity, we write $\hat{\pi}=\hat{\pi}(f)$ when the choice of $f$ is clear from the context.
More generally, in our analysis in Section 3, we need to calculate the total weight within a subset ${\mathcal{E}}\subseteq\binom{S}{2}$ given by
$\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})\triangleq\sum_{e\in{\mathcal{E}}}\beta_{e}(\mathcal{H}_{\pi}^{f}).$ (7)
Due to the correlation between the edges in $G_{1}$ and $G_{2}$, the summands $\beta_{e}(\mathcal{H}_{\pi}^{f})$ are correlated random variables. The main idea is to decompose ${\mathcal{E}}$ into independent parts. Specially, the correlation is governed by the underlying mapping $\pi^{*}$, as illustrated in Figure 1 where the correlated edges are represented by red dashed lines. To formally describe all correlation relationships, we introduce the correlated functional digraph of a mapping $\pi$ between a pair of graphs.
[…]
Figure 1: Examples of the mapping $\pi$ and the underlying correlation $\pi^{*}$, where the domain and range of $\pi$ and $\pi^{*}$ could be different.
Definition 6 (Correlated functional digraph).
Let $\pi^{*}:S^{*}\mapsto T^{*}$ be the underlying mapping between correlated elements. The correlated functional digraph of the function $\pi:S\mapsto T$ is constructed as follows. Let the vertex set be $\binom{S}{2}\cup\binom{S^{*}}{2}\cup\binom{T}{2}\cup\binom{T^{*}}{2}$. We first add every edge $e\mapsto\pi(e)$ for $e\in\binom{S}{2}$, and then merge each pair of nodes $(e,\pi^{*}(e))$ for $e\in\binom{S^{*}}{2}$ into one node.
[…]
Figure 2: The connected components in the correlated functional digraph.
It should be noted that both $\pi$ and $\pi^{*}$ are injective mappings under our model. After merging all pairs of nodes according to $\pi^{*}$, the degree of each vertex in the correlated functional digraph is at most two. Therefore, the connected components consist of paths and cycles, where a self-loop is understood as a cycle of length one. The connected components are illustrated in Figure 2. Let ${\mathcal{P}}$ and ${\mathcal{C}}$ denote the collections of subsets of ${\mathcal{E}}$ belonging to different connected paths and cycles, respectively. Note that the sets from ${\mathcal{P}}$ and ${\mathcal{C}}$ are disjoint. Consequently,
$\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})=\sum_{P\in{\mathcal{P}}}\beta_{P}(\mathcal{H}_{\pi}^{f})+\sum_{C\in{\mathcal{C}}}\beta_{C}(\mathcal{H}_{\pi}^{f}),$
where the summands are mutually independent.
In our models, edge correlations are assumed to be homogeneous, implying that the distributions of $\beta_{P}(\mathcal{H}_{\pi}^{f})$ and $\beta_{C}(\mathcal{H}_{\pi}^{f})$ depend only on the size of the component. Let $\kappa^{{\mathsf{P}},f}_{\ell}(t)$ and $\kappa^{{\mathsf{C}},f}_{\ell}(t)$ denote the cumulant generating functions of $\beta_{P}(\mathcal{H}_{\pi}^{f})$ and $\beta_{C}(\mathcal{H}_{\pi}^{f})$ for components of order $\ell$, respectively. Then, we have
$\log\mathbb{E}\left[e^{t\beta_{P}(\mathcal{H}_{\pi}^{f})}\right]=\kappa^{{\mathsf{P}},f}_{|P|}(t),\qquad\log\mathbb{E}\left[e^{t\beta_{C}(\mathcal{H}_{\pi}^{f})}\right]=\kappa^{{\mathsf{C}},f}_{|C|}(t).$
For simplicity, we write $\kappa_{\ell}^{\mathsf{C}}(t)=\kappa_{\ell}^{{\mathsf{C}},f}(t)$ and $\kappa_{\ell}^{\mathsf{P}}(t)=\kappa_{\ell}^{{\mathsf{P}},f}(t)$ when the function $f$ is specified. Lower-order cumulants can be calculated directly. See more details in Appendix B.1. However, it is crucial to establish upper bounds for higher-order cumulants in terms of the lower-order ones. To this end, we introduce the following lemma.
Lemma 1.
For any $10<p,\rho<$ 1,
$\kappa^{{\mathsf{P}}}_{1}(t)\leq\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\leq\kappa^{{\mathsf{C}}}_{1}(t)\quad\text{and}\quad\kappa^{{\mathsf{P}}}_{\ell}(t)\leq\kappa^{{\mathsf{C}}}_{\ell}(t)\leq\frac{\ell}{2}\kappa^{{\mathsf{C}}}_{2}(t),\quad\forall~{}\ell\geq 2,$
under any of the following three conditions:
  • In the Erdős-Rényi model with $f(x,y)=xy$ and $t>0$;
  • In the Gaussian Wigner model with $f(x,y)=xy$ and $0<t<\frac{1}{1+\rho}$;
  • In the Gaussian Wigner model with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ and $t>0$.
Consequently,
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})}\right]\leq\frac{|{\mathcal{E}}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$ (8)
where $L$ denotes the number of self-loops.
The proof of Lemma 1 is deferred to Appendix B.1. The special case where both $\pi$ and $\pi^{*}$ are bijective has been studied in [WXY22, DD23b, HM23], where the correlation relationships can be characterized by the permutation $(\pi^{*})^{-1}\circ\pi$. In this case, the connected components of the functional digraph of permutations are all cycles. In contrast, in our setting, the domains and ranges of $\pi$ and $\pi^{*}$ may differ, requiring us to deal with the intricate correlations among edges involving both cycles and paths, as addressed by Lemma 1.
Remark 2.
For the Gaussian Wigner model, the MLE is defined by (6) with the following similarity function:
$f(x,y)=-\frac{\rho^{2}}{2}\left(x^{2}+y^{2}\right)+\rho xy.$ (9)
We analyze approximations of the above similarity score under different regimes. Specifically, when $\rho=1-\Omega(1)$, we apply the estimator with $f(x,y)=xy$, whereas for $\rho=1-o(1)$ the estimator with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ turns out to be crucial.
3 Recovery by maximizing total similarity score
In this section, we establish the possibility results by analyzing the estimators $\hat{\pi}$ defined in (6). For any two injections $\pi:S\mapsto T$ and $\pi^{\prime}:S^{\prime}\mapsto T^{\prime}$ with $|S|=|S^{\prime}|=m$, let $d(\pi,\pi^{\prime})\triangleq m-|\{v\in S\cap S^{\prime}:\pi(v)=\pi^{\prime}(v)\}|$. Indeed, $d(\pi,\pi^{\prime})=m(1-\mathrm{overlap}(\pi,\pi^{\prime}))$, where $\mathrm{overlap}(\pi,\pi^{\prime})$ is defined in (3). By the optimality condition, it suffices to show that
$\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})>\max_{\pi:d(\pi,\pi^{*})\geq d_{0}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})=\max_{k\geq d_{0}}\max_{\pi:d(\pi,\pi^{*})=k}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})$
with high probability, where the thresholds $d_{0}=1$ and $(1-\delta)m$ correspond to exact and partial recoveries, respectively. In the following, we outline a general recipe to derive an upper bound for the failure event $\{\max_{\pi\in\mathcal{T}_{k}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})\}$ for a fixed $k$, where $\mathcal{T}_{k}\subseteq\mathcal{S}_{n,m}$ denotes the set of injections $\pi$ such that $d(\pi,\pi^{*})=k$. The error probabilities in the main theorems are then obtained by summing over the corresponding range of $k$.
For any $\pi\in\mathcal{T}_{k}$, by definition, there exists a set of correctly matched vertices of cardinality $m-k$—corresponding to the self-loops in the correlated functional digraph of $\pi$ over the vertices—denoted by $F_{\pi}\triangleq\{v\in S^{*}\cap S:\pi^{*}(v)=\pi(v)\}$ with $|F_{\pi}|=m-k$. The induced subgraphs of ${\mathcal{H}}_{\pi}^{f}$ and ${\mathcal{H}}_{\pi^{*}}^{f}$ over $F_{\pi}$ are identical. Consequently,
$\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})\iff\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}]).$
It should be noted that correlated random variables are present on both sides of the inequality. Nevertheless, for any threshold $\tau_{k}$, either $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}$ or $\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}$ must hold. This leads to the following upper bound:
$\left\{\max_{\pi\in\mathcal{T}_{k}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})\right\}\subseteq\bigcup_{\pi\in{\mathcal{T}}_{k}}\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\}\cup\{\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}\}.$ (10)
The first event indicates the presence of a weak signal, while the second reflects the impact of strong noise. The key result to establish is that, for a suitable $\tau_{k}$, the probabilities of these bad events are small. Here, $\tau_{k}$ can be selected as a function of $m,k,p,\rho$, and $f$. For brevity, we write $\tau_{k}=\tau(m,k,p,\rho,f)$.
Bad event of signal.
For a fixed $\pi\in\mathcal{T}_{k}$, the random variable $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])$ represents the total weight across $N_{k}\triangleq\binom{m}{2}-\binom{m-k}{2}=mk(1-\frac{k+1}{2m})$ pairs of vertices. Additionally, $F_{\pi}$ is a subset of $S^{*}$ of cardinality $m-k$, and the total number of possible configurations for $F_{\pi}$ is at most $\binom{m}{m-k}=\binom{m}{k}$. Therefore,
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\mathbb{P}\left[\bigcup_{\begin{subarray}{c}F\subseteq S^{*}\\ |F|=m-k\end{subarray}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F])<\tau_{k}\right\}\right]$
$\leq\binom{m}{k}\mathbb{P}\left[\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])<\tau_{k}\right].$ (11)
Bad event of noise.
The analysis of the noise component is more challenging due to the mismatch between $\pi$ and the underlying $\pi^{*}$. Let $S_{\pi}$ denote the domain of $\pi$, and ${\mathcal{E}}_{\pi}\triangleq\binom{S_{\pi}}{2}-\binom{F_{\pi}}{2}$ with $|{\mathcal{E}}_{\pi}|=N_{k}$. The total weight $\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])$ can be equivalently expressed as $\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})$. The cumulant generating function for this quantity is upper bounded in Lemma 1, utilizing the decomposition provided by the correlated functional digraph. As a result, the error probability can be evaluated using the Chernoff bound by optimizing over $t>0$ in (8).
To this end, we need to upper bound the number of self-loops in (8). For a self-loop over an edge $e=uv$, we have $\pi(uv)=\pi^{*}(uv)$. Note that ${\mathcal{E}}_{\pi}$ excludes the edges in the induced subgraph over $F_{\pi}$. Therefore, it necessarily holds that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$, which contribute two mismatched vertices in the reconstruction of the underlying mapping. Since the total number of mismatched vertices for $\pi\in{\mathcal{T}}_{k}$ equals $k$, the number of self-loops is at most $\frac{k}{2}$. Consequently, applying (8) with the formula for lower-order cumulants provides an upper bound for $\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]$.
It remains to upper bound the total number of $\pi$ in $\mathcal{T}_{k}$. To do so, we first choose $m-k$ elements from the domain of $\pi^{*}$ and map them to the same value as $\pi^{*}$. Then, the remaining domain and range, each of size $k$, are matched arbitrarily. This yields:
$|\mathcal{T}_{k}|\leq\binom{m}{m-k}\binom{n-m+k}{k}^{2}k!\overset{\mathrm{(a)}}{\leq}\frac{m^{k}n^{2k}}{k!^{2}}\overset{\mathrm{(b)}}{\leq}\frac{n^{3k}}{k!^{2}},$
where $\mathrm{(a)}$ uses the bound $\binom{m}{m-k}\leq\frac{m^{k}}{k!}$ and $\binom{n-m+k}{k}\leq\frac{(n-m+k)^{k}}{k!}\leq\frac{n^{k}}{k!}$, and $\mathrm{(b)}$ is because $m\leq n$. Therefore,
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}\right\}\right]\leq\frac{n^{3k}}{k!^{2}}\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right].$ (12)
3.1 Erdős-Rényi model
In this subsection, we focus on the Erdős-Rényi model and use $\hat{\pi}$ from (6) with $f(x,y)=xy$ as the estimator. For the bad event of signal, since $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])\sim\mathrm{Bin}(N_{k},p_{11})$ with $|F|=k$, the error probability follows from the standard Chernoff bound (see, e.g., (52)): for $10<\eta<$ 1,
$\mathbb{P}\left[\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right]\leq\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right),\quad\tau_{k}=N_{k}p_{11}(1-\eta).$ (13)
For the bad event of noise, applying (8) along with the lower-order cumulant (36) and (37) yields the following lemma, whose proof is deferred to Appendix B.2.
Lemma 2.
In the Erdős-Rényi model, for $f(x,y)=xy$ and $\pi\in\mathcal{S}_{n,m}$ with $k=d(\pi,\pi^{*})$, if $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$, then
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}+\frac{k\gamma}{4(2+\gamma)}\right).$ (14)
Since $p_{11}=p^{2}(1+\gamma)$, we need to choose an appropriate $\eta\in(0,1)$ depending on the model parameters, such that $(1+\gamma)(1-\eta)>1$, in order to apply the error probabilities in (13) and (14) to upper bound (11) and (12), respectively. When $m$ exceeds the sample complexities for partial and exact recovery, we explicitly construct $\eta$ to prove the following propositions, which establish the possibility results stated in Theorems 1 and 2. The proofs of Propositions 1 and 2 are deferred to Appendix A.1 and A.2, respectively.
Proposition 1 (Erdős-Rényi model, upper bound for partial recovery).
For any $\delta\in(0,1)$, there exists a constant $c_{1}(\delta)>0$ such that, when $m\geq\frac{c_{1}(\delta)\log n}{p^{2}\phi(\gamma)}$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=xy$ satisfies
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\leq\frac{4\log m+2}{m}.$
Proposition 2 (Erdős-Rényi model, upper bound for exact recovery).
There exists a universal constant $C_{1}>0$ such that, when $m\geq C_{1}\left(\frac{\log(1/p^{2}\gamma)}{p^{2}\gamma}\vee\frac{\log n}{p^{2}\phi(\gamma)}\right)$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=xy$ satisfies
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\leq\frac{1}{m-1}+\frac{1}{n-1}.$
3.2 Gaussian Wigner model
In this subsection, we focus on the Gaussian Wigner model. To establish upper bounds for exact recovery, we use different estimators depending on the value of $\rho$, where $0<\rho\leq 1-e^{-12}$ corresponds to the weak signal regime, and $11-e^{-12}<\rho<$ 1 corresponds to the strong signal regime.
Weak signal
Under the weak signal regime $0<\rho\leq 1-e^{-12}$, we use $\hat{\pi}$ in (6) with $f(x,y)=xy$ as our estimator. Similar to Erdős-Rényi model, we upper bound the event $\{d(\hat{\pi},\pi^{*})=k\}$ by the bad event of signal and the bad event of noise according to (10). For the bad event of signal, we note that $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])=\sum_{i=1}^{N_{k}}A_{i}B_{i}$, where $(A_{i},B_{i})$ are independent and identically distributed (i.i.d.) pairs of standard normals with correlation coefficient $\rho$. Then, the error probability can be upper bounded by
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\binom{m}{k}\mathbb{P}\left[\sum_{i=1}^{N_{k}}A_{i}B_{i}<\tau_{k}\right],$ (15)
the tail follows from the Hanson-Wright inequality in Lemma 9. For the bad event of noise, applying (8) with the formula of lower-order cumulants (41) and (42) yields the following Lemma, whose proof is deferred to Appendix B.3.
Lemma 3.
In the Gaussian Wigner model, for $f(x,y)=xy$, any threshold $\tau_{k}$, and $\pi\in\mathcal{S}_{n,m}$ with $k=d(\pi,\pi^{*})$, we have
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)\geq\tau_{k}\right]\leq\exp\left(-\frac{\rho\tau_{k}}{6}+\frac{\rho^{2}|\mathcal{E}_{\pi}|}{14}+\frac{\log 5}{8}k\right).$ (16)
The following proposition provides sufficient condition on $m$ for exact recovery in the Gaussian Wigner model when $0<\rho\leq 1-e^{-12}$, whose proof is deferred to Appendix A.3.
Proposition 3 (Gaussian Wigner model, upper bound for exact recovery).
When $0<\rho\leq 1-e^{-12}$, there exists a universal constant $C_{3}>0$ such that, when $m\geq\frac{C_{3}\log n}{\rho^{2}}$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=xy$ satisfies
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\leq\frac{1}{m-1}+\frac{1}{n-1}.$
Strong signal
Under the strong signal regime $11-e^{-12}<\rho<$ 1, we use $\hat{\pi}$ in (6) with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ as our estimator. Indeed, the MLE is also of the form (6), but with $f(x,y)=\rho xy-\frac{\rho^{2}}{2}(x^{2}+y^{2})$, which is approximately $-\frac{1}{2}(x-y)^{2}$ when $\rho=1-o(1)$. This choice reflects the phase transition phenomenon in the Gaussian Wigner model when $\rho=1-o(1)$. Specifically, for a correlated edge pair, the expected mean squared difference $\mathbb{E}\left[\left(\beta_{e}(G_{1})-\beta_{\pi^{*}(e)}(G_{2})\right)^{2}\right]=o(1)$, while for an uncorrelated edge pair, it stays bounded away from zero with high probability. Our proof follows a similar structure with the weak signal regime. For the bad event of signal, since $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])=\sum_{i=1}^{N_{k}}-\frac{1}{2}(A_{i}-B_{i})^{2}$, where $(A_{i},B_{i})$ are i.i.d. pairs of standard normals with correlation coefficient $\rho$. Then, the error probability can be upper bounded by
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\binom{m}{k}\mathbb{P}\left[\sum_{i=1}^{N_{k}}-\frac{1}{2}(A_{i}-B_{i})^{2}<\tau_{k}\right],$ (17)
the tail follows from the concentration inequality for chi-squared distribution in Lemma 10. For the bad event of noise, applying (8) with the formula of lower-order cumulants (45) and (46) yields the following Lemma, whose proof is deferred to Appendix B.4.
Lemma 4.
In the Gaussian Wigner model, for $f(x,y)=-\frac{1}{2}(x-y)^{2}$, any threshold $\tau_{k}$, and $\pi\in\mathcal{S}_{n,m}$ with $k=d(\pi,\pi^{*})$, we have
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right)+\frac{k}{8}\log\left(\frac{1}{1-\rho}\right)\right).$
The following proposition provides a sufficient condition on $m$ for exact recovery in the Gaussian Wigner model when $11-e^{-12}<\rho<$ 1, whose proof is deferred to Appendix A.4.
Proposition 4.
When $11-e^{-12}<\rho<$ 1, there exists a universal constant $C_{4}>0$ such that, when $m\geq C_{4}\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right)$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ satisfies
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\leq\frac{2}{n-1}.$
In view of Propositions 3 and 4, we note that $\rho^{2}\asymp\log\left(\frac{1}{1-\rho^{2}}\right)$ when $0<\rho\leq 1-e^{-12}$, and $\log\left(\frac{1}{1-\rho}\right)\asymp\log\left(\frac{1}{1-\rho^{2}}\right)$ when $11-e^{-12}<\rho<$ 1. Then, there exists a universal constant $C_{2}$ such that,
$C_{2}\left(\frac{\log n}{\log(1/(1-\rho^{2}))}\vee 1\right)\geq\begin{cases}\frac{C_{3}\log n}{\rho^{2}},&0<\rho\leq 1-e^{-12},\\ C_{4}\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right),&1-e^{-12}<\rho<1.\end{cases}$
Therefore, we prove the possibility results in Theorem 3.
Remark 3.
While the proofs for the possibility results in the Erdős-Rényi and Gaussian Wigner models share a similar structure, they also exhibit several key differences. Following the intuition of MLE in (9), we analyze two different estimators for the Gaussian Wigner model. Furthermore, by applying different concentration inequalities, we observe a significant difference in the choice of $\tau_{k}$. Specifically, we set $\tau_{k}=\Theta\left(\mathbb{E}\left[\mathsf{{e}}\left(\mathcal{H}_{\pi^{*}}^{f}\right)-\mathsf{{e}}\left(\mathcal{H}_{\pi^{*}}^{f}[F_{\pi}]\right)\right]\right)$ for both models with $\rho=1-\Omega(1)$. In contrast, we choose $\tau_{k}=\omega\left(\mathbb{E}\left[\mathsf{{e}}\left(\mathcal{H}_{\pi^{*}}^{f}\right)-\mathsf{{e}}\left(\mathcal{H}_{\pi^{*}}^{f}[F_{\pi}]\right)\right]\right)$ for the Gaussian model with $\rho=1-o(1)$. The thresholds are chosen to balance the error probabilities arising from the bad events of signal and noise, ultimately achieving the optimal rates for the correlated size $m$.
Remark 4.
Our estimator (6) requires knowledge of the cardinality $|S^{*}|=m$. When $m$ is unknown, the objective function in (6) will increase monotonically with $m$. To address this, we can introduce a penalty term for $m$ in the estimator to identify the correct size, following an approach similar to the MLE in the new model. This will be discussed in more detail in a forthcoming working paper.
4 Impossibility results
In this section, we present the impossibility results for the graph alignment problem. Under our proposed model, the alignment problem aims to recover the domain $S^{*}\subseteq V(G_{1})$, range $T^{*}\subseteq V(G_{2})$, and the mapping $\pi^{*}:S^{*}\mapsto T^{*}$. When equipped with the additional knowledge on $S^{*}$ and $T^{*}$, our problem can be reduced to recovery with full observations on smaller graphs, the reconstruction threshold for which is settled in [WXY22]. The lower bound therein remains valid when the number of correlated nodes is substituted with $m$. However, such reduction only proves tight in a limited number of regimes (see Proposition 6). We will establish the impossibility results for the remaining regimes by Fano’s method. Two main ingredients of Fano’s method are outlined as follows:
  • Construct a packing set ${\mathcal{M}}$ of the parameter space $\mathcal{S}_{n,m}$ such that any two distinct elements from ${\mathcal{M}}$ differ by a prescribed threshold. Specifically, in partial recovery, the overlap of each pair is less than $\delta$, which is equivalent to $\min_{\pi\neq\pi^{\prime}\in{\mathcal{M}}}d(\pi,\pi^{\prime})\geq(1-\delta)m$, while in exact recovery ${\mathcal{M}}=\mathcal{S}_{n,m}$. The cardinality of ${\mathcal{M}}$ measures the complexity of the parameter space under the target metric.
  • Choose the uniform prior on $\pi^{*}$ over ${\mathcal{M}}$ and upper bound the mutual information $I(\pi^{*};G_{1},G_{2})$. Given $\pi^{*}$, the conditional distribution of the observed graphs $(G_{1},G_{2})$ is specified in Definition 4. For the mutual information, let $\mathcal{P}$ denote the joint distribution of $(G_{1},G_{2})$ and $\mathcal{Q}$ be any distribution over $(G_{1},G_{2})$. Then,
    $I(\pi^{*};G_{1},G_{2})=\mathbb{E}_{\pi^{*}}[D(\mathcal{P}_{G_{1},G_{2}|\pi^{*}}\|\mathcal{P}_{G_{1},G_{2}})]\leq\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}}).$ (18)
    The impossibility results follow if $I(\pi^{*};G_{1},G_{2})\leq c\log|{\mathcal{M}}|$ for some small constant $c$.
Let ${\mathcal{M}}_{\delta}$ denote a packing set under the overlap threshold $\delta$. The size of ${\mathcal{M}}_{\delta}$ follows from the standard volume argument [PW25, Theorem 27.3]. For $r\in[m]$, let $B(\pi,r)\triangleq\{\pi^{\prime}:d(\pi,\pi^{\prime})\leq r\}$ denote the ball of radius $r$ centered at $\pi$. Then, we obtain
$|{\mathcal{M}}_{\delta}|\geq\frac{|\mathcal{S}_{n,m}|}{\max_{\pi}|B(\pi,(1-\delta)m-1)|}\geq\frac{|\mathcal{S}_{n,m}|}{\max_{\pi}|B(\pi,(1-\delta)m)|}.$
It remains to evaluate the cardinality of $\mathcal{S}_{n,m}$ and upper bound the volume of the ball under our distance metric $d$. It is straightforward to obtain that $|\mathcal{S}_{n,m}|=\binom{n}{m}^{2}m!.$ Let $k=\delta m$. Note that all elements from $B(\pi,m-k)$ have at least $k$ common mappings. To upper bound $|B(\pi,m-k)|$, we first choose $k$ elements from the domain of $\pi$ and map to the same value as $\pi$, and the remaining domain and range of size $m-k$ and the mapping are selected arbitrarily. We get $|B(\pi,m-k)|\leq\binom{m}{k}\binom{n-k}{m-k}^{2}(m-k)!.$ Consequently,
$|{\mathcal{M}}_{\delta}|\geq\frac{\binom{n}{m}^{2}m!}{\binom{m}{k}\binom{n-k}{m-k}^{2}(m-k)!}=\left(\frac{\binom{n}{k}}{\binom{m}{k}}\right)^{2}k!>\left(\frac{n^{2}k}{e^{3}m^{2}}\right)^{k}\geq\left(\frac{\delta n}{e^{3}}\right)^{k},$ (19)
where we use the inequalities that $(\frac{n}{k})^{k}\leq\binom{n}{k}<(\frac{en}{k})^{k}$ and $k!\geq(k/e)^{k}$. Fano’s method provides a lower bound on the Bayesian risk when $\pi$ is uniformly distributed over ${\mathcal{M}}_{\delta}$, which further lower bounds the minimax risk. The above argument also yields a lower bound when $\pi$ is uniform over $\mathcal{S}_{n,m}$ via generalized Fano’s inequality [BHM12, Lemma 20]. For the Erdős-Rényi model, the following propositions provide lower bounds for $m$ for partial recovery and exact recovery, and thus prove the lower bounds in Theorems 1 and 2. For the Gaussian Wigner model, the Proposition 5 provides a lower bound for $m$ for partial recovery, which implies that $m\leq\frac{c\log n}{\log\left(1/(1-\rho^{2})\right)}\vee\frac{1}{2}$ is sufficient for the lower bound since $m<1$ is impossible for partial recovery. Consequently, we finish the proof of impossibility results in Theorem 3 since $\frac{c\log n}{\log\left(1/(1-\rho^{2})\right)}\vee\frac{1}{2}\asymp\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1$.
Proposition 5 (Lower bound for partial recovery).
In the Erdős-Rényi model, for any $\delta\in(0,1)$, if $m\leq\frac{c\log n}{p^{2}\phi(\gamma)}$, then for any estimator $\hat{\pi}$,
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\geq 1-\frac{13c}{\delta}.$
In the Gaussian Wigner model, for any $\delta\ in(0,1)$, if $m\leq\frac{c\log n}{\log(1/(1-\rho^{2}))}$, then for any estimator $\hat{\pi}$,
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\geq 1-\frac{c}{2\delta}.$
Proof.
In the Erdős-Rényi model, for any $\pi$ with domain $S$ and range $T$ such that $|S|=|T|=m$, we arbitrarily pick a bijection $\sigma:V(G_{1})\mapsto V(G_{2})$ such that $\sigma|_{S}=\pi$. Then, the conditional distribution $\mathcal{P}_{G_{1},G_{2}|\pi}$ can be factorized into
$\mathcal{P}_{G_{1},G_{2}|\pi}=\prod_{e\in\binom{S}{2}}P(e,\pi(e))\prod_{e\in\binom{V(G_{1})}{2}\backslash\binom{S}{2}}Q(e,\sigma(e)),$
where $P=\mathrm{Bern}(p,p,\rho)$ and $Q=\mathrm{Bern}(p,p,0)$. Pick $\mathcal{Q}$ in (18) to be an auxiliary null model under which $G_{1}$ and $G_{2}$ are independent with the same marginal as $\mathcal{P}$. Then, $\mathcal{Q}_{G_{1},G_{2}}$ can be factorized into
$\mathcal{Q}_{G_{1},G_{2}}=\prod_{e\in\binom{S}{2}}Q(e,\pi(e))\prod_{e\in\binom{V(G_{1})}{2}\backslash\binom{S}{2}}Q(e,\sigma(e)).$
The KL-divergence between the product measures $\mathcal{P}_{G_{1},G_{2}|\pi}$ and $\mathcal{Q}_{G_{1},G_{2}}$ can be expressed as
$D\left(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}}\right)=\binom{m}{2}D(P\|Q)$
for any $\pi:S\mapsto T$ with $|S|=|T|=m$. Applying Lemma 7, we obtain
$\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq 25\binom{m}{2}p^{2}\phi(\gamma).$ (20)
Applying generalized Fano’s inequality [BHM12, Lemma 20] with (19) and (20), we obtain
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{25\binom{m}{2}p^{2}\phi(\gamma)}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{13c}{\delta},$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
In the Gaussian Wigner model, let $P=\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ and $Q=\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right)$. We can similarly pick the auxiliary null model $\mathcal{Q}$ under which $G_{1}$ and $G_{2}$ are independent. Then, $\mathcal{Q}$ can be factorized into
$\mathcal{Q}_{G_{1},G_{2}}=\prod_{e\in\binom{S}{2}}Q(e,\pi(e))\prod_{e\in\binom{V(G_{1})}{2}\backslash\binom{S}{2}}Q(e,\sigma(e)).$
Applying Lemma 7, we obtain
$\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right).$ (21)
Applying generalized Fano’s inequality [BHM12, Lemma 20] with (19) and (21), we obtain
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right)}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{c}{2\delta},$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$. ∎
Proposition 6 (Erdős-Rényi model, lower bound for exact recovery).
For any $c\in(0,1)$ and any estimator $\hat{\pi}$, there exists a constant $c_{3}$ depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)\right)$,
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\geq 1-c,$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
Proof.
We first apply the reduction argument. With the additional information on the domain and range of $\pi^{*}$, our problem can be reduced to the reconstruction of mapping as in [WXY22]. Applying the lower bound in [WXY22, Theorem 4], for a fixed $\epsilon\in(0,1)$, when $m(\sqrt{p_{00}p_{11}}-\sqrt{p_{01}p_{10}})^{2}\leq(1-\epsilon)\log m$, we have $\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\geq 1-o(1)$ for any estimator $\hat{\pi}$. Note that $(\sqrt{p_{00}p_{11}}-\sqrt{p_{01}p_{10}})^{2}\asymp p^{2}(\gamma\wedge\gamma^{2})\asymp(\rho^{2})\wedge(\rho p).$ Therefore, when
$m\lesssim\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\log\left(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\right),$ (22)
we have $\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\geq 1-o(1)$. Applying Proposition 5 with $\delta=1/2$ yields that, when
$m\lesssim\frac{\log n}{p^{2}\phi(\gamma)},$ (23)
we have $\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\geq 1-c$ for $c\in(0,1)$.
When $\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\asymp n$, by (22), exact recovery is impossible, even when $m=n$. Next, we consider the regime that $\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\lesssim n$. When $\gamma\leq 1$, we have $p^{2}(\gamma\wedge\gamma^{2})=p^{2}\gamma^{2}\asymp p^{2}\phi(\gamma)$, and thus
$\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\log\left(\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\right)\lesssim\frac{\log n}{p^{2}\phi(\gamma)}.$
When $\gamma\geq 1$, $\gamma\wedge\gamma^{2}=\gamma$. By comparing (22) and (23), we derive that exact recovery is impossible if $m\lesssim\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)$. ∎
5 Discussion and future directions
This paper introduces the partially correlated Erdős-Rényi model and the partially correlated Gaussian Wigner model, wherein a pair of induced subgraphs of a specific size is correlated. We investigate the optimal information-theoretic threshold for recovering both the latent correlated subgraphs and the hidden vertex correspondence in these new models. In comparison with prior work on the correlated Erdős-Rényi and correlated Gaussian Wigner models, the additional challenge arises from the unknown location of the correlated subsets. For a candidate mapping $\pi$ whose domain may include both correlated and ambient vertices, we extend the classical notion of functional digraph to formally describe the correlation structure among the edges. From the correlated functional digraph, we observe that the independent components consist of cycles and paths. The graphical representation may be of independent interest for general models.
There are many problems to be further investigated under our proposed models:
  • Refined results. The results in the paper could be further refined in various ways, such as deriving the sharp constants and characterizing the optimal scaling in terms of the fraction $\delta$ in partial recovery.
  • Efficient algorithms. It is of interest to investigate the polynomial-time algorithms and identify the computational hardness under our model. More efficient algorithms are also desirable when the signal is stronger.
  • Graph sampling. One motivation of the paper stems from graph sampling as discussed in Section 1.2. The sampled subgraphs are partially correlated, where the size of correlated subsets is a random variable depending on the sampling methods. Thus, it is natural to ask about the sample size needed for reliable recovery.
  • Correlation test. The correlation test problem under our model is also highly relevant. It is interesting to find out whether the detection problem is strictly easier than recovery, both in terms of the information thresholds and algorithmic developments.
Appendix A Proof of Propositions
A.1 Proof of Proposition 1
For any $k\geq(1-\delta)m$, let $\tau_{k}=N_{k}p_{11}(1-\eta)$ with
$\eta=\sqrt{\frac{8h(\frac{k}{m})}{kp_{11}}}\cdot{\mathbf{1}_{\left\{{k\leq m-1}\right\}}}+\sqrt{\frac{\log n}{kmp_{11}}}\cdot{\mathbf{1}_{\left\{{k=m}\right\}}},$
where $h(x)\triangleq-x\log x-(1-x)\log(1-x)$ is the binary entropy function and $p_{11}=p^{2}(1+\gamma)$. Let $c_{1}(\delta)=100\vee\frac{200h(1-\delta)}{1-\delta}$. We first show that $\eta\leq\frac{\gamma}{4(1+\gamma)}<1$ under $m\geq\left(100\vee\frac{200h(1-\delta)}{1-\delta}\right)\frac{\log n}{p^{2}\phi(\gamma)}$. Since $\phi(\gamma)=(1+\gamma)\log(1+\gamma)-\gamma\leq\gamma^{2}$ by $\log(1+\gamma)\leq\gamma$, we obtain that $p^{2}\phi(\gamma)\leq p^{2}\gamma^{2}=\rho^{2}(1-p)^{2}\leq 1$ and $m\geq c_{1}(\delta)\log n$. We note that
$\eta$ $\overset{\mathrm{(a)}}{\leq}\sqrt{\frac{8h(1-\delta)}{(1-\delta)mp_{11}}}\cdot{\mathbf{1}_{\left\{{k\leq m-1}\right\}}}+\sqrt{\frac{\log n}{m^{2}p_{11}}}\cdot{\mathbf{1}_{\left\{{k=m}\right\}}}$
$\overset{\mathrm{(b)}}{\leq}\left(\sqrt{\frac{8h(1-\delta)}{1-\delta}}\vee\frac{1}{\sqrt{c_{1}(\delta)}}\right)\frac{1}{\sqrt{mp_{11}}}$
$\overset{\mathrm{(c)}}{\leq}\left(\sqrt{\frac{8h(1-\delta)}{(1-\delta)c_{1}(\delta)}}\vee\frac{1}{c_{1}(\delta)}\right)\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}}\overset{\mathrm{(d)}}{\leq}\frac{1}{5}\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}},$
where $\mathrm{(a)}$ uses the fact that $\frac{h(k/m)}{k/m}\leq\frac{h(1-\delta)}{1-\delta}$ since $\frac{h(x)}{x}$ decreases in $(0,1)$ and $k\geq(1-\delta)m$; $\mathrm{(b)}$ follows from $m\geq c_{1}(\delta)\log n$; $\mathrm{(c)}$ is because $mp_{11}=mp^{2}(1+\gamma)\geq\frac{c_{1}(\delta)\log n}{\phi(\gamma)/(1+\gamma)}=\frac{c_{1}(\delta)\log n}{\log(1+\gamma)-\gamma/(1+\gamma)}$; $\mathrm{(d)}$ follows from $c_{1}(\delta)=100\vee\frac{200h(1-\delta)}{1-\delta}$. Recall the assumption stated in Section 1.1, where it’s asserted that $p\geq n^{-1}$ in the Erdős-Rényi model, thereby implying $\log(1+\gamma)\leq\log n$ and thus $\eta\leq\frac{1}{5}$. When $\gamma>10$, $\eta\leq\frac{1}{5}\leq\frac{\gamma}{4(1+\gamma)}$. When $\gamma\leq 10$, since $\log(1+x)-\frac{x}{1+x}-x^{2}\leq 0$ for any $x>0$, we obtain $\eta\leq\frac{1}{5}\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}}\leq\frac{\gamma}{5\sqrt{\log n}}\leq\frac{\gamma}{4(1+\gamma)}$ when $n$ sufficiently large. Therefore, we obtain $\eta\leq\frac{\gamma}{4(1+\gamma)}<1$.
We then upper bound the bad event of signal by Chernoff bound. By [CT06, Lemma 17.5.1], we have $\binom{m}{k}\leq\exp\left[mh(k/m)\right]$. When $k\leq m-1$, it follows from (11) and (13) that
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\binom{m}{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right)$
$\leq\exp\left(mh\left(\frac{k}{m}\right)-\frac{N_{k}p_{11}\eta^{2}}{2}\right)$
$\leq\exp\left(-mh\left(\frac{k}{m}\right)\right),$ (24)
where the last inequality follows from $\frac{N_{k}p_{11}\eta^{2}}{2}=2m\left(2-\frac{k+1}{m}\right)h\left(\frac{k}{m}\right)\geq 2mh\left(\frac{k}{m}\right)$ when $k\leq m-1$. When $k=m$, $N_{k}=\frac{mk}{2}\left(2-\frac{k+1}{m}\right)\geq\frac{mk}{3}$. Then, it follows from (11) and (13) that
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]\leq\binom{m}{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right)\leq n^{-1/6},$ (25)
where the last inequality follows from $\binom{m}{k}=1$ and $\frac{N_{k}p_{11}\eta^{2}}{2}=\frac{\log n}{2}\frac{N_{k}}{mk}\geq\frac{\log n}{6}$ when $k=m$.
Next, we upper bound the bad event of noise. Since $\eta\leq\frac{\gamma}{4(1+\gamma)}$, it follows from Lemma 5 that $\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}=(1+\gamma)(1-\eta)>1$, where $p_{11}=p^{2}(1+\gamma)$ and $|\mathcal{E}_{\pi}|=N_{k}$. Since $e^{\frac{k\gamma}{4(2+\gamma)}}\leq e^{\frac{k}{4}}$ and $k!\geq\left(\frac{k}{e}\right)^{k}$, it follows from (12) and Lemma 2 that
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}\right\}\right]$ $\leq n^{3k}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right)$
$=n^{3k}\exp\left(-\frac{N_{k}p^{2}}{2}\phi\left((1+\gamma)(1-\eta)-1\right)\right)$
$\overset{\mathrm{(a)}}{\leq}n^{3k}\exp\left(-\frac{N_{k}p^{2}\phi(\gamma)}{8}\right)\overset{\mathrm{(b)}}{\leq}n^{-k},$ (26)
where $\mathrm{(a)}$ applies (49) in Lemma 5; $\mathrm{(b)}$ is because $N_{k}=\frac{mk}{2}\left(2-\frac{k+1}{m}\right)\geq\frac{mk}{3}$ and $\frac{mp^{2}\phi(\gamma)}{24}\geq 4\log n$.
Finally, we upper bound the error probability $\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]$. By summing (24) over $(1-\delta)m\leq k\leq m-1$ and summing (26) over $(1-\delta)m\leq k\leq m$ and (25), we obtain that
$~{}\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]$
$=$ $~{}\sum_{k=(1-\delta)m}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]$
$\leq$ $~{}\sum_{k=(1-\delta)m}^{m}\left\{\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]+\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}\right\}\right]\right\}$
$\leq$ $~{}n^{-1/6}+\sum_{k=(1-\delta)m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]+\sum_{k=(1-\delta)m}^{m}n^{-k}$
$\leq$ $~{}n^{-1/6}+\sum_{k=(1-\delta)m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]+\frac{n^{-(1-\delta)m}}{1-n^{-1}}.$
Since $m\geq c_{1}(\delta)\log n$ and $c_{1}(\delta)\geq 100$, it follows from Lemma 6 that $\sum_{k=(1-\delta)m}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]\leq\frac{4\log m+2}{m}.$
A.2 Proof of Proposition 2
Let $\tau_{k}=N_{k}p_{11}(1-\eta)$ with $\eta=\frac{\gamma}{4(1+\gamma)}<1$. By (11) and (13) and applying $\binom{m}{k}\leq m^{k}$, we get
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]\leq m^{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right).$ (27)
Since $p_{11}=p^{2}(1+\gamma)$ and $|\mathcal{E}_{\pi}|=N_{k}$, we have $\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}=(1+\gamma)(1-\eta)>1$ by Lemma 5. Since $e^{\frac{k\gamma}{4(2+\gamma)}}\leq e^{\frac{k}{4}}$ and $k!\geq\left(\frac{k}{e}\right)^{k}$, we obtain $\frac{1}{k!^{2}}e^{\frac{k\gamma}{4(2+\gamma)}}\leq 1$. Then it follows from (12) and Lemma 2 that
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}\right\}\right]$ $\leq n^{3k}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right)$
$=n^{3k}\exp\left(-\frac{N_{k}p^{2}}{2}\phi\left((1+\gamma)(1-\eta)-1\right)\right)$
$\leq n^{3k}\exp\left(-\frac{N_{k}p^{2}}{8}\phi(\gamma)\right),$ (28)
where the last inequality applies (49) in Lemma 5. By summing (27) and (28) over $k\geq 1$ and applying $N_{k}\geq\frac{km}{3}$, we obtain that
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]=$ $~{}\sum_{k=1}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]$
$\leq$ $~{}\sum_{k=1}^{m}\left\{\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]+\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi}^{f}[F_{\pi}])\geq\tau_{k}\right\}\right]\right\}$
$\leq$ $~{}\sum_{k=1}^{m}\left\{\left[m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\right]^{k}+\left[n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\right]^{k}\right\}.$ (29)
Let $C_{1}=3000$. It remains to upper bound (29) under $m\geq 3000\left(\frac{\log(1/p^{2}\gamma)}{p^{2}\gamma}\vee\frac{\log n}{p^{2}\phi(\gamma)}\right)$. Since $m\geq\frac{3000\log n}{p^{2}\phi(\gamma)}$, we have
$n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\leq\frac{1}{n}.$ (30)
With a weak signal that $\gamma\leq 1$, we have $\phi(\gamma)\leq\frac{\gamma^{2}}{1+\gamma}$, and thus $mp_{11}\eta^{2}=\frac{mp^{2}\gamma^{2}}{16(1+\gamma)}\geq\frac{mp^{2}\phi(\gamma)}{16}$. Since $m\geq\frac{3000\log n}{p^{2}\phi(\gamma)}$, we have
$m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\leq\frac{1}{m}.$ (31)
With a strong signal that $\gamma\geq 1$, we have $\frac{\gamma}{1+\gamma}\geq\frac{1}{2}$ and thus $mp_{11}\eta^{2}=\frac{mp^{2}\gamma^{2}}{16(1+\gamma)}\geq\frac{mp^{2}\gamma}{32}$. Since $m\geq 3000\left(\frac{1}{p^{2}\gamma}\log\frac{1}{p^{2}\gamma}\right)$, we have that
$\frac{\log m}{m}$ $\overset{\mathrm{(a)}}{\leq}\frac{{\log 3000+\log\left(\frac{1}{p^{2}\gamma}\right)+\log\left(\log\left(\frac{1}{p^{2}\gamma}\right)\right)}}{3000\left(\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)\right)}\overset{\mathrm{(b)}}{\leq}\frac{\left(2+\frac{\log 3000}{\log 4}\right)\log\left(\frac{1}{p^{2}\gamma}\right)}{3000\left(\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)\right)}\leq\frac{p^{2}\gamma}{384},$
where $\mathrm{(a)}$ is because $\frac{\log x}{x}$ decreases on $[e,\infty]$ and $m\geq 3000$; $\mathrm{(b)}$ is because $\log\left(\frac{1}{p^{2}\gamma}\right)=\log\left(\frac{1}{\rho p(1-p)}\right)\geq\log 4$ and $\log\left(\log\left(\frac{1}{p^{2}\gamma}\right)\right)\leq\log\left(\frac{1}{p^{2}\gamma}\right)$. Then (31) holds since $m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\leq m\exp\left(-\frac{mp^{2}\gamma}{192}\right)\leq m\exp(-2\log m)=m^{-1}$. We conclude the proof by applying (29) with (30) and (31).
A.3 Proof of Proposition 3
Let $\tau_{k}=\rho N_{k}-c_{0}\left(\sqrt{N_{k}\log(1/\theta)}\vee\log(1/\theta)\right)$ with $\theta=\exp\left(-2k\log m\right)$, where $c_{0}$ is the universal constant in Lemma 9. Let $C_{3}=25c_{0}^{2}\vee 1100$. We first verify that $\tau_{k}\geq\frac{1}{2}\rho N_{k}$ under condition $m\rho^{2}\geq C_{3}\log n$. We show that $c_{0}\sqrt{N_{k}\log(1/\theta)}\leq\frac{\rho N_{k}}{2}$ and $c_{0}\log(1/\theta)\leq\frac{\rho N_{k}}{2}$, respectively. For the first term $c_{0}\sqrt{N_{k}\log(1/\theta)}$, we note that
$c_{0}\sqrt{N_{k}\log(1/\theta)}$ $\overset{\mathrm{(a)}}{\leq}\sqrt{\frac{2C_{3}}{25}N_{k}k\log n}=\sqrt{\frac{C_{3}\log n}{m}\cdot\frac{mk}{3}\cdot\frac{6N_{k}}{25}}\overset{\mathrm{(b)}}{\leq}\sqrt{\rho^{2}\cdot N_{k}\cdot\frac{6N_{k}}{25}}\leq\frac{1}{2}\rho N_{k},$
where $\mathrm{(}a)$ follows from $25c_{0}^{2}\leq C_{3}$ and $\log(1/\theta)=2k\log m\leq 2k\log n$; $\mathrm{(b)}$ follows from $m\rho^{2}\geq C_{3}\log n$ and $N_{k}=mk\left(1-\frac{k+1}{2m}\right)\geq\frac{mk}{3}$. For the second term $c_{0}\log(1/\theta)$, we have that
$c_{0}\log(1/\theta)$ $\overset{\mathrm{(a)}}{\leq}6c_{0}N_{k}\frac{\log m}{m}\overset{\mathrm{(b)}}{\leq}\frac{6c_{0}}{\sqrt{C_{3}}}N_{k}\cdot\sqrt{\frac{C_{3}\log n}{m}}\cdot\sqrt{\frac{\log m}{m}}\overset{\mathrm{(c)}}{\leq}\frac{6c_{0}\sqrt{\log m}}{\sqrt{C_{3}m}}\rho N_{k}\leq\frac{1}{2}\rho N_{k}$
when $m$ sufficiently large, where $\mathrm{(a)}$ is because $\log(1/\theta)=2k\log m$ and $N_{k}\geq\frac{mk}{3}$; $\mathrm{(b)}$ follows from $\log m\leq\log n$; $\mathrm{(c)}$ uses $m\rho^{2}\geq C_{3}\log n$. Therefore, we obtain $\tau_{k}\geq\frac{1}{2}\rho N_{k}$.
For the bad event of signal, by (15) and Hanson-Wright inequality in Lemma 9 with $M_{0}=I_{N_{k}}$ and applying $\binom{m}{k}\leq m^{k}$, we obtain that
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]\leq m^{k}\exp\left(-2k\log m\right)=m^{-k}.$ (32)
For the bad event of noise, it follows from (12) and Lemma 3 that
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $~{}n^{3k}\exp\left(-\frac{\rho\tau_{k}}{6}+\frac{\rho^{2}|\mathcal{E}_{\pi}|}{14}+\frac{\log 5}{8}k\right)$
$\overset{\mathrm{(a)}}{\leq}$ $~{}n^{3k}\exp\left(-\frac{\rho^{2}|\mathcal{E}_{\pi}|}{84}+\frac{\log 5}{8}k\right)\overset{\mathrm{(b)}}{\leq}n^{3k}\exp\left(-4k\log n\right)=n^{-k},$ (33)
where $\mathrm{(a)}$ is because $\tau_{k}\geq\frac{1}{2}\rho N_{k}$; $\mathrm{(b)}$ follows from $|\mathcal{E}_{\pi}|=N_{k}\geq\frac{mk}{3}$ and $m\rho^{2}\geq 1100\log n$. By summing (32) and (A.3) over $1\leq k\leq m$, we obtain that
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]=$ $~{}\sum_{k=1}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]$
$\leq$ $~{}\sum_{k=1}^{m}\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]+\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $~{}\sum_{k=1}^{m}\left(m^{-k}+n^{-k}\right)\leq\frac{m^{-1}}{1-m^{-1}}+\frac{n^{-1}}{1-n^{-1}}=\frac{1}{m-1}+\frac{1}{n-1}.$
A.4 Proof of Proposition 4
Let $\tau_{k}=\frac{\rho-1}{3}\log\left(\frac{1}{1-\rho}\right)N_{k}$ and $C_{4}=100$. We first upper bound the bad event of signal. Recall (17), since $A_{i}-B_{i}\sim\mathcal{N}(0,2-2\rho)$ for $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$, we obtain that $\sum_{i=1}^{N_{k}}\frac{(A_{i}-B_{i})^{2}}{2(1-\rho)}$ follows the chi-squared distribution with $N_{k}$ degrees of freedom. Then, it follows from (17) and $\binom{m}{k}\leq m^{k}$ that
$~{}\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]$
$\leq$ $~{}m^{k}\mathbb{P}\left[\sum_{i=1}^{N_{k}}-\frac{1}{2}(A_{i}-B_{i})^{2}<\frac{\rho-1}{3}\log\left(\frac{1}{1-\rho}\right)N_{k}\right]$
$=$ $~{}m^{k}\mathbb{P}\left[\sum_{i=1}^{N_{k}}\frac{(A_{i}-B_{i})^{2}}{2(1-\rho)}>\frac{1}{3}\log\left(\frac{1}{1-\rho}\right)N_{k}\right]$
$\overset{\mathrm{(a)}}{\leq}$ $~{}m^{k}\exp\left\{-\frac{N_{k}}{2}\left[\frac{1}{3}\log\left(\frac{1}{1-\rho}\right)-1-\log\left(\frac{1}{3}\log\left(\frac{1}{1-\rho}\right)\right)\right]\right\}$
$\overset{\mathrm{(b)}}{\leq}$ $~{}m^{k}\exp\left(-\frac{mk\log\left(1/(1-\rho)\right)}{48}\right)\overset{\mathrm{(c)}}{\leq}n^{-k},$ (34)
where $\mathrm{(a)}$ uses (53) in Lemma 10; $\mathrm{(b)}$ is because $N_{k}=\binom{m}{2}-\binom{m-k}{2}=mk\left(1-\frac{k+1}{2m}\right)\geq\frac{km}{3}$ and $x-1-\log x\geq\frac{3x}{8}$ for $x=\frac{1}{3}\log\left(\frac{1}{1-\rho}\right)\geq 4$; $\mathrm{(c)}$ is because $m\log(1/(1-\rho))\geq 100\log n$ and $n\geq m$.
We then upper bound the bad event of noise. It follows from (12) and Lemma 4 that
$~{}\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $~{}n^{3k}\exp\left(-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right)+\frac{k}{8}\log\left(\frac{1}{1-\rho}\right)\right)$
$\overset{\mathrm{(a)}}{\leq}$ $~{}n^{3k}\exp\left(-\frac{|\mathcal{E}_{\pi}|}{6}\log\left(\frac{1}{1-\rho}\right)+\frac{k}{8}\log\left(\frac{1}{1-\rho}\right)\right)$
$\overset{\mathrm{(b)}}{\leq}$ $~{}\exp\left(3k\log n-\frac{mk}{18}\log\left(\frac{1}{1-\rho}\right)+\frac{k}{8}\log\left(\frac{1}{1-\rho}\right)\right)$
$=$ $~{}\exp\left[\left(3k\log n-\frac{mk}{24}\log\frac{1}{1-\rho}\right)+\left(\frac{k}{8}\log\frac{1}{1-\rho}-\frac{mk}{72}\log\frac{1}{1-\rho}\right)\right]\overset{\mathrm{(c)}}{\leq}n^{-k},$ (35)
where $\mathrm{(a)}$ is because $|\mathcal{E}_{\pi}|=N_{k}$ and $-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right)={\frac{(\rho-3)|\mathcal{E}_{\pi}|}{12}}\log\left(\frac{1}{1-\rho}\right)\leq-\frac{|\mathcal{E}_{\pi}|}{6}\log\left(\frac{1}{1-\rho}\right)$; $\mathrm{(b)}$ follows from $|\mathcal{E}_{\pi}|=N_{k}=\frac{mk}{2}\left(1-\frac{k+1}{2m}\right)\geq\frac{mk}{3}$; $\mathrm{(c)}$ is because $m\geq 100\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right)$ implies $3k\log n-\frac{mk}{24}\log\frac{1}{1-\rho}\leq-k\log n$ and $\frac{k}{8}\log\frac{1}{1-\rho}-\frac{mk}{72}\log\frac{1}{1-\rho}\leq 0$. By summing over (34) and (35), we obtain that
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]=$ $~{}\sum_{k=1}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]$
$\leq$ $~{}\sum_{k=1}^{m}\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]+\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}^{f}_{\pi})-\mathsf{{e}}({\mathcal{H}}^{f}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $~{}\sum_{k=1}^{m}\left(n^{-k}+n^{-k}\right)\leq\frac{2n^{-1}}{1-n^{-1}}=\frac{2}{n-1}.$
Appendix B Proof of Lemmas
B.1 Proof of Lemma 1
Erdős-Rényi Model
We consider $f(x,y)=xy$ in the Erdős-Rényi model. The lower-order cumulants can be calculated directly:
$\kappa^{{\mathsf{C}}}_{1}(t)$ $=\log(1+p_{11}(e^{t}-1)),\quad\kappa^{{\mathsf{P}}}_{1}(t)=\log(1+p^{2}(e^{t}-1)),$ (36)
$\kappa^{{\mathsf{C}}}_{2}(t)$ $=\log(1+2p^{2}(e^{t}-1)+p_{11}^{2}(e^{t}-1)^{2}).$ (37)
We first evaluate the moment generating function for paths. Consider a path $P$ of size $\ell$ denoted by $\langle e_{1}e_{2}\dots e_{\ell}\rangle$ as illustrated in Figure 3. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$. By definition (7),
$\beta_{P}(\mathcal{H}_{\pi}^{f})=\sum_{i=1}^{\ell}\beta_{e_{i}}(G_{1})\beta_{\pi(e_{i})}(G_{2})=\sum_{i=1}^{\ell}A_{i-1}B_{i}.$
[…]
Figure 3: Illustration of a path of size $\ell$.
For notational simplicity, we introduce an auxiliary random variable $B_{0}$ that is correlated with $A_{0}$ such that $(A_{0},B_{0})\sim\mathrm{Bern}(p,p,\rho)$. Then
$m_{\ell}\triangleq\mathbb{E}\left[e^{t\beta_{P}\left(\mathcal{H}^{f}_{\pi}\right)}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\prod_{i=1}^{\ell}e^{tA_{i-1}B_{i}}|B_{0}\dots B_{\ell}\right]\right]=\mathbb{E}\left[\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}B_{i}}|B_{i-1},B_{i}\right]\right]$
$=\sum_{b_{0},\dots,b_{\ell}\in\{0,1\}}\prod_{i=0}^{\ell}\mathbb{P}\left[B_{i}=b_{i}\right]\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right].$ (38)
Define $M(b_{i-1},b_{i})\triangleq\mathbb{P}\left[B_{i}=b_{i}\right]\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right]$ for $b_{i-1},b_{i}\in\{0,1\}$ and a matrix
$M\triangleq\begin{bmatrix}M(0,0)&M(0,1)\\ M(1,0)&M(1,1)\end{bmatrix}=\begin{bmatrix}\bar{p}&(\bar{p}+p_{01}(e^{t}-1))p/\bar{p}\\ \bar{p}&p+p_{11}(e^{t}-1)\end{bmatrix},$
where $\bar{p}\triangleq 1-p$. Recall that $\mathbb{P}\left[B_{i}=1\right]=p$. Then, we obtain that
$m_{\ell}=\sum_{b_{0},\dots,b_{\ell}\in\{0,1\}}\mathbb{P}\left[B_{0}=b_{0}\right]M(b_{0},b_{1})\dots M(b_{\ell-1},b_{\ell})=[\bar{p},p]M^{\ell}\begin{bmatrix}1\\ 1\end{bmatrix}.$
The trace and determinant of $M$ are given by
$T\triangleq\mathrm{Tr}(M)=1+p_{11}(e^{t}-1),\quad D\triangleq\det(M)=\rho p\bar{p}(e^{t}-1)>0.$
Since $D<p_{11}(e^{t}-1)$, the discriminant is $T^{2}-4D>0$. Hence, the matrix $M$ has two distinct eigenvalues denoted by $\lambda_{1}>\lambda_{2}$. Since $\lambda_{1}+\lambda_{2}=T>0$ and $\lambda_{1}\lambda_{2}=D>0$, we have $\lambda_{1}>\lambda_{2}>0$, and the general term of $m_{\ell}$ is
$m_{\ell}=\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}.$ (39)
The coefficients $\alpha_{1}$ and $\alpha_{2}$ can be determined via the first two terms $m_{0}=1$ and $m_{1}$. Then we get
$m_{\ell}=\left(\frac{1}{2}+\frac{2m_{1}-T}{2\sqrt{T^{2}-4D}}\right)\lambda_{1}^{\ell}+\left(\frac{1}{2}-\frac{2m_{1}-T}{2\sqrt{T^{2}-4D}}\right)\lambda_{2}^{\ell}.$
Furthermore, by plugging $m_{1}=1+p^{2}(e^{t}-1)$, we get $T-m_{1}=D$ and thus $m_{1}(T-m_{1})>D$, which is equivalent to $|2m_{1}-T|<\sqrt{T^{2}-4D}$. Therefore, both coefficients $\alpha_{1},\alpha_{2}\in(0,1)$.
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$ for $i=0,\dots,\ell-1$.
[…]
Figure 4: Illustration of a cycle of size $\ell$.
Following a similar argument as (38), we have
$\tilde{m}_{\ell}\triangleq\mathbb{E}\left[e^{t\beta_{C}\left(\mathcal{H}^{f}_{\pi}\right)}\right]$ $=\sum_{b_{1},\dots,b_{\ell}=b_{0}\in\{0,1\}}\prod_{i=1}^{\ell}\mathbb{P}\left[B_{i}=b_{i}\right]\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right]$
$=\sum_{b_{1},\dots,b_{\ell}=b_{0}\in\{0,1\}}M(b_{0},b_{1})M(b_{1},b_{2})\dots M(b_{\ell-1},b_{0}).$
Applying the eigenvalue decomposition of $M$ again, we obtain that
$\tilde{m}_{\ell}=\mathrm{Tr}(M^{\ell})=\lambda_{1}^{\ell}+\lambda_{2}^{\ell}.$ (40)
By definition, $\kappa^{{\mathsf{P}}}_{\ell}(t)=\log m_{\ell}$ and $\kappa^{{\mathsf{C}}}_{\ell}(t)=\log\tilde{m}_{\ell}$. To upper bound the cumulants, it suffices to consider $m_{\ell}$ and $\tilde{m}_{\ell}$. In (39), we have $\alpha_{1},\alpha_{2}\in(0,1)$ and $\lambda_{1}>\lambda_{2}>0$. By monotonicity, it follows that $m_{\ell}\leq\tilde{m}_{\ell}$ and thus
$\kappa^{{\mathsf{P}}}_{\ell}(t)\leq\kappa^{{\mathsf{C}}}_{\ell}(t).$
For $x\in\mathbb{R}^{n}$ and $\ell\geq 2$, we have $\|{x}\|_{\ell}\leq\|{x}\|_{2}\leq\|{x}\|_{1}$. It follows from the formula of $\tilde{m}_{\ell}$ in (40) that $\tilde{m}_{\ell}^{1/\ell}\leq\tilde{m}_{2}^{1/2}\leq\tilde{m}_{1}$. Equivalently,
$\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\leq\kappa^{{\mathsf{C}}}_{1}(t),\quad\kappa^{{\mathsf{C}}}_{\ell}(t)\leq\frac{\ell}{2}\kappa^{{\mathsf{C}}}_{2}(t)\quad\forall~{}\ell\geq 2.$
The last inequality $2\kappa^{{\mathsf{P}}}_{1}(t)\leq\kappa^{{\mathsf{C}}}_{2}(t)$ follows by comparing the explicit formula $\kappa^{{\mathsf{P}}}_{1}(t)=\log(1+p^{2}(e^{t}-1))$ with $\kappa^{{\mathsf{C}}}_{2}(t)$ in (37) and using $p_{11}\geq p^{2}$.
Finally, since the summands over different connected components are independent, it follows that
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})}\right]$ $=\sum_{P\in{\mathcal{P}}}\kappa_{|P|}^{{\mathsf{P}}}(t)+\sum_{C\in{\mathcal{C}}}\kappa_{|C|}^{{\mathsf{C}}}(t)$
$\leq\sum_{P\in{\mathcal{P}}}\frac{|P|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+\sum_{C\in{\mathcal{C}}:|C|\geq 2}\frac{|C|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+\sum_{C\in{\mathcal{C}}:|C|=1}\kappa_{1}^{{\mathsf{C}}}(t)$
$=\frac{|{\mathcal{E}}|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+|\{C\in{\mathcal{C}}:|C|=1\}|\left(\kappa_{1}^{{\mathsf{C}}}(t)-\frac{1}{2}\kappa_{2}^{{\mathsf{C}}}(t)\right),$
where the last equality uses fact that $|{\mathcal{E}}|=\sum_{P\in{\mathcal{P}}}|P|+\sum_{C\in{\mathcal{C}}}|C|$.
Remark 5.
We have two bounds for large $\ell$ in the Erdős-Rényi model, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)\leq\frac{1}{\ell}\log(\lambda_{1}^{\ell}+\lambda_{2}^{\ell})$, where $10<\alpha_{2}<\alpha_{1}<1,\alpha_{1}+\alpha_{2}=$ 1 and $\lambda_{1}>\lambda_{2}>0$. Consequently, $\lambda_{1}-\frac{\log 2}{\ell}\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq\lambda_{1}+\frac{\log 2}{\ell}.$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $\|x\|_{\ell}\leq\|x\|_{2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
Gaussian Wigner Model, Part 1
In this part, we focus on the Gaussian Wigner model with $f(x,y)=xy$. The lower-order cumulants can be calculated directly:
$\kappa_{1}^{\mathsf{C}}(t)$ $=-\frac{1}{2}\log(1-2t\rho-t^{2}(1-\rho^{2})),\quad\kappa_{1}^{\mathsf{P}}(t)=-\frac{1}{2}\log\left(1-t^{2}\right),$ (41)
$\kappa_{2}^{\mathsf{C}}(t)$ $=-\frac{1}{2}\log\left(1-2t^{2}(1+\rho^{2})+t^{4}(1-\rho^{2})^{2}\right).$ (42)
We first evaluate the moment generating function for paths. Consider a path $P$ of size $\ell$ denoted by $\langle e_{1}e_{2}\dots e_{\ell}\rangle$ as illustrated in Figure 3. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$. By definition (7),
$\beta_{P}(\mathcal{H}^{f}_{\pi})=\sum_{i=1}^{\ell}\beta_{e_{i}}(G_{1})\beta_{\pi(e_{i})}(G_{2})=\sum_{i=1}^{\ell}A_{i-1}B_{i}.$
For the sake of notational simplicity, we introduce an auxiliary variable $B_{0}$ that is correlated with $A_{0}$ such that $(A_{0},B_{0})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$. Then
$m_{\ell}\triangleq\mathbb{E}\left[e^{t\beta_{P}\left(\mathcal{H}^{f}_{\pi}\right)}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\prod_{i=0}^{\ell-1}e^{tA_{i}B_{i+1}}|B_{0}\dots B_{\ell}\right]\right]=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\mathbb{E}\left[e^{tA_{i}B_{i+1}}|B_{i},B_{i+1}\right]\right]$
$=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\exp\left(t\rho B_{i}B_{i+1}+\frac{1}{2}t^{2}(1-\rho^{2})B_{i+1}^{2}\right)\right],$ (43)
where the last equality follows from $A_{i}|B_{i}\sim\mathcal{N}(\rho B_{i},1-\rho^{2})$ and $\mathbb{E}\left[\exp(tZ)\right]=\exp\left(t\mu+t^{2}\nu^{2}/2\right)$ for $Z\sim\mathcal{N}(\mu,\nu^{2})$.
Define $\mathbf{B}_{\ell}\triangleq[B_{0},B_{1},\cdots,B_{\ell}]$ and a matrix
$\mathbf{W}_{\ell}\triangleq\begin{bmatrix}1&-t\rho&0&\cdots&0\\ -t\rho&1-t^{2}(1-\rho^{2})&-t\rho&\cdots&0\\ 0&-t\rho&1-t^{2}(1-\rho^{2})&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&-t\rho\\ 0&0&\cdots 0&-t\rho&1-t^{2}(1-\rho^{2})\end{bmatrix},$
where $(\mathbf{W}_{\ell})_{00}=1,(\mathbf{W}_{\ell})_{ii}=1-t^{2}(1-\rho^{2})$ for any $1\leq i\leq\ell$ and $(\mathbf{W}_{\ell})_{ij}=-t\rho$ for any $|i-j|=1$ with $0\leq i,j\leq\ell$. By (43), we have
$m_{\ell}$ $=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\exp\left(t\rho B_{i}B_{i+1}+\frac{1}{2}t^{2}(1-\rho^{2})B_{i+1}^{2}\right)\right]$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell}B_{i}^{2}\right)\exp\left(\sum_{i=0}^{\ell-1}\left(t\rho B_{i}B_{i+1}+\frac{1}{2}t^{2}(1-\rho^{2})B_{i+1}^{2}\right)\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\mathbf{B}_{\ell}^{\top}\mathbf{W}_{\ell}\mathbf{B}_{\ell}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}.$ (44)
If $\mathbf{W}_{\ell}$ is positive definite, then $m_{\ell}=\det\left(\mathbf{W}_{\ell}\right)^{-1/2}$ by (44). We then prove $\mathbf{W}_{\ell}$ is positive definite and compute the explicit formula of $\det\left(\mathbf{W}_{\ell}\right)$. Expanding the last column of $\mathbf{W}_{\ell}$ yields that
$\det(\mathbf{W}_{\ell})=\left(1-t^{2}(1-\rho^{2})\right)\det(\mathbf{W}_{\ell-1})-t^{2}\rho^{2}\det(\mathbf{W}_{\ell-2}),\text{ for any }\ell\geq 2.$
Therefore, the general term of $\det(\mathbf{W}_{\ell})$ is determined by
$\det(\mathbf{W}_{\ell})=\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell},$
where $\lambda_{1}>\lambda_{2}$ are two roots of equation $x^{2}-\left(1-t^{2}(1-\rho^{2})\right)x+t^{2}\rho^{2}=0$. The coefficients $\alpha_{1}$ and $\alpha_{2}$ can be determined via the first two terms $\det(\mathbf{W}_{0})=1$ and $\det(\mathbf{W}_{1})=1-t^{2}$. Specifically,
$\alpha_{1}={\frac{1}{2}+\frac{1-t^{2}(1+\rho^{2})}{2\sqrt{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}}}},\quad\alpha_{2}={\frac{1}{2}-\frac{1-t^{2}(1+\rho^{2})}{2\sqrt{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}}}}.$
Since $01-t^{2}(1-\rho^{2})-2t\rho=(1-t\rho-t)(1-t\rho+t)>$ 0 by $t<\frac{1}{1+\rho}$, we obtain $(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}>0$, and hence
$\frac{1-t^{2}(1+\rho^{2})}{\sqrt{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}}}=\sqrt{\frac{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}+4t^{4}\rho^{2}}{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}}}>1,$
which implies that $\alpha_{1}>1$ and $\alpha_{2}<0$. Since $\lambda_{1}+\lambda_{2}=1-t^{2}(1-\rho^{2})>0$ by $t<\frac{1}{1+\rho}$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}>0$ and the discriminant $(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}>0$ , we have $\lambda_{1}>\lambda_{2}>0$. Therefore,
$\det(\mathbf{W}_{\ell})=\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}=\alpha_{1}(\lambda_{1}^{\ell}-\lambda_{2}^{\ell})+\lambda_{2}^{\ell}>0,\text{ for any }\ell\geq 0,$
which implies that $\mathbf{W}_{\ell}$ is positive definite by Sylvester’s criterion (see, e.g., [HJ12, Theorem 7.2.5]). Denote $\mathbf{W}_{\ell}=\mathbf{J}_{\ell}^{\top}\mathbf{J}_{\ell}$ with $\mathbf{J}_{\ell}\in\mathbb{R}^{(\ell+1)\times(\ell+1)}$. Then, by (44),
$m_{\ell}$ $=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\mathbf{B}_{\ell}^{\top}\mathbf{W}_{\ell}\mathbf{B}_{\ell}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\mathbf{B}_{\ell}^{\top}\mathbf{J}_{\ell}^{\top}\mathbf{J}_{\ell}\mathbf{B}_{\ell}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}=\left[\det\left(\mathbf{J}_{\ell}\right)\right]^{-1}=\left[\det\left(\mathbf{W}_{\ell}\right)\right]^{-1/2}.$
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for $i=0,\dots,\ell-1$. Recall that $\lambda_{1}+\lambda_{2}=1-t^{2}(1-\rho^{2})$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}$. Following a similar argument as (43), we have
$\tilde{m}_{\ell}$ $\triangleq\mathbb{E}\left[e^{t\beta_{C}(\mathcal{H}^{f}_{\pi})}\right]=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\exp\left(t\rho B_{i}B_{i+1}+\frac{1}{2}t^{2}(1-\rho^{2})B_{i+1}^{2}\right)\right]$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell-1}B_{i}^{2}\right)\exp\left(\sum_{i=0}^{\ell-1}\left(t\rho B_{i}B_{i+1}+\frac{1}{2}t^{2}(1-\rho^{2})B_{i+1}^{2}\right)\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell-1}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell-1}\left(\lambda_{1}^{1/2}B_{i}-\lambda_{2}^{1/2}B_{i+1}\right)^{2}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell-1}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell}\exp\left(-\frac{1}{2}\mathbf{B}_{\ell-1}^{\top}\tilde{\mathbf{J}}_{\ell-1}^{\top}\tilde{\mathbf{J}}_{\ell-1}\mathbf{B}_{\ell-1}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell-1}=\left[\det(\tilde{\mathbf{J}}_{\ell-1})\right]^{-1},$
where
$\tilde{\mathbf{J}}_{\ell-1}\triangleq\begin{bmatrix}\lambda_{1}^{1/2}&-\lambda_{2}^{1/2}&0&\cdots&0\\ 0&\lambda_{1}^{1/2}&-\lambda_{2}^{1/2}&\cdots&0\\ 0&0&\lambda_{1}^{1/2}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&-\lambda_{2}^{1/2}\\ -\lambda_{2}^{1/2}&0&\cdots 0&0&\lambda_{1}^{1/2}\end{bmatrix}$
and hence $\det(\tilde{\mathbf{J}}_{\ell-1})=\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}$. Therefore,
$m_{\ell}^{-2}-\tilde{m}_{\ell}^{-2}$ $=\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}-\left(\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}\right)^{2}$
$=(\alpha_{1}-1)\left(\lambda_{1}^{\ell}-\lambda_{2}^{\ell}\right)+\lambda_{2}^{\ell/2}\left(\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}\right)+\lambda_{1}^{\ell/2}\lambda_{2}^{\ell/2}>0,$
which implies that $m_{\ell}<\tilde{m}_{\ell}$ for any $\ell\geq 1$.
Next, we prove that $\tilde{m}_{\ell}\leq\left(\tilde{m}_{2}\right)^{\ell/2}$ for $\ell\geq 2$. Indeed, since $\left(\frac{\lambda_{1}-\lambda_{2}}{\lambda_{1}}\right)^{\ell/2}+\left(\frac{\lambda_{2}}{\lambda_{1}}\right)^{\ell/2}\leq\frac{\lambda_{1}-\lambda_{2}}{\lambda_{1}}+\frac{\lambda_{2}}{\lambda_{1}}=1$ when $\ell\geq 2$, we obtain $\tilde{m}_{\ell}^{-1}=\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}\geq(\lambda_{1}-\lambda_{2})^{\ell/2}=\tilde{m}_{2}^{-\ell/2}$. We also have $\tilde{m}_{2}^{-1}-\tilde{m}_{1}^{-2}=2\lambda_{2}^{1/2}\left(\lambda_{1}^{1/2}-\lambda_{2}^{1/2}\right)>0$. Recall that $\kappa_{\ell}^{\mathsf{C}}(t)=\log\left(\tilde{m}_{\ell}\right),\kappa_{\ell}^{\mathsf{C}}(t)=\log\left(m_{\ell}\right)$, and $m_{\ell}<\tilde{m}_{\ell}$. Consequently,
$\frac{1}{2}\kappa_{2}^{\mathsf{C}}(t)\leq\kappa_{1}^{\mathsf{C}}(t),\quad\kappa_{\ell}^{{\mathsf{P}}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)\quad\forall\ell\geq 2.$
The last inequality $\kappa_{1}^{\mathsf{P}}(t)\leq\frac{1}{2}\kappa_{2}^{\mathsf{C}}(t)$ follows from
$m_{1}^{-4}-\tilde{m}_{2}^{-2}=\left(1-t^{2}\right)^{2}-\left[1-2t^{2}-2t^{2}\rho^{2}+t^{4}(1-\rho^{2})^{2}\right]=t^{2}\rho^{2}\left(2+2t^{2}-t^{2}\rho^{2}\right)\overset{\mathrm{(a)}}{>}0,$
where $\mathrm{(a)}$ is because $t^{2}\rho^{2}<\frac{\rho^{2}}{(1+\rho)^{2}}<1$ and $12+2t^{2}>$ 1.
Finally, since the summands over different connected components are independent, it follows that
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})}\right]$ $=\sum_{P\in{\mathcal{P}}}\kappa_{|P|}^{{\mathsf{P}}}(t)+\sum_{C\in{\mathcal{C}}}\kappa_{|C|}^{{\mathsf{C}}}(t)$
$\leq\sum_{P\in{\mathcal{P}}}\frac{|P|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+\sum_{C\in{\mathcal{C}}:|C|\geq 2}\frac{|C|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+\sum_{C\in{\mathcal{C}}:|C|=1}\kappa_{1}^{{\mathsf{C}}}(t)$
$=\frac{|{\mathcal{E}}|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+|\{C\in{\mathcal{C}}:|C|=1\}|\left(\kappa_{1}^{{\mathsf{C}}}(t)-\frac{1}{2}\kappa_{2}^{{\mathsf{C}}}(t)\right),$
where the last equality uses fact that $|{\mathcal{E}}|=\sum_{P\in{\mathcal{P}}}|P|+\sum_{C\in{\mathcal{C}}}|C|$.
Gaussian Wigner Model, Part 2
In this part, we focus on the Gaussian Wigner model with $f(x,y)=-\frac{1}{2}(x-y)^{2}$. The lower-order cumulants can be calculated directly:
$\kappa_{1}^{\mathsf{C}}(t)$ $=-\frac{1}{2}\log\left(1+2t(1-\rho)\right),\quad\kappa_{1}^{\mathsf{P}}(t)=-\frac{1}{2}\log(1+2t),$ (45)
$\kappa_{2}^{\mathsf{C}}(t)$ $=-\frac{1}{2}\log\left((1+2t)^{2}-4t^{2}\rho^{2}\right).$ (46)
We first evaluate the moment generating function for paths. Consider a path $P$ of size $\ell$ denoted by $\langle e_{1}e_{2}\dots e_{\ell}\rangle$ as illustrated in Figure 3. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$. By definition (7),
$\beta_{P}(\mathcal{H}_{\pi}^{f})=\sum_{i=1}^{\ell}-\frac{1}{2}\left(\beta_{e_{i}}(G_{1})-\beta_{\pi(e_{i})}(G_{2})\right)^{2}=\sum_{i=1}^{\ell}-\frac{1}{2}\left(A_{i-1}-B_{i}\right)^{2}.$
For the sake of notational simplicity, we introduce an auxiliary variable $B_{0}$ that is correlated with $A_{0}$ such that $(A_{0},B_{0})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$. Then
$m_{\ell}\triangleq\mathbb{E}\left[e^{t\beta_{P}\left(\mathcal{H}_{\pi}^{f}\right)}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\prod_{i=0}^{\ell-1}e^{-\frac{t}{2}\left(A_{i}-B_{i+1}\right)^{2}}|B_{0}\dots B_{\ell}\right]\right]=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\mathbb{E}\left[e^{-\frac{t}{2}\left(A_{i}-B_{i+1}\right)^{2}}|B_{i},B_{i+1}\right]\right]$
$=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\frac{1}{\sqrt{1+t(1-\rho^{2})}}\exp\left(\frac{-\frac{t}{2}\left(\rho B_{i}-B_{i+1}\right)^{2}}{1+t(1-\rho^{2})}\right)\right],$ (47)
where the last equation follows from $(A_{i}-B_{i+1})\mid(B_{i},B_{i+1})\sim\mathcal{N}(\rho B_{i}-B_{i+1},1-\rho^{2})$ and $\mathbb{E}\left[\exp(-tZ^{2})\right]=\frac{1}{\sqrt{1+2t\nu^{2}}}\exp\left(\frac{-\mu^{2}t}{1+2t\nu^{2}}\right)$ for $Z\sim\mathcal{N}(\mu,\nu^{2})$ when $t>0$.
Define $\mathbf{B}_{\ell}\triangleq[B_{0},B_{1},\cdots,B_{\ell}]$ and a matrix
$\mathbf{W}_{\ell}\triangleq\begin{bmatrix}1+t&-t\rho&0&\cdots&0\\ -t\rho&1+2t&-t\rho&\cdots&0\\ 0&-t\rho&1+2t&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&-t\rho\\ 0&0&\cdots 0&-t\rho&1+t(2-\rho^{2})\end{bmatrix},$
where $(\mathbf{W}_{\ell})_{00}=1+t,(\mathbf{W}_{\ell})_{\ell\ell}=1+t(2-\rho^{2}),(\mathbf{W}_{\ell})_{ii}=1+2t$ for any $11\leq i\leq\ell-$ 1 and $(\mathbf{W}_{\ell})_{ij}=-t\rho$ for any $|i-j|=1$ with $0\leq i,j\leq\ell$. We note that the symmetric matrix $\mathbf{W}_{\ell}$ is strictly diagonally dominant with positive diagonal entries, and then it is positive definite. This follows from the eigenvalues of symmetric matrix being real, and the Gershgorin’s circle theorem [Ger31]. Denote $\mathbf{W}_{\ell}=\mathbf{J}_{\ell}^{\top}\mathbf{J}_{\ell}$ with $\mathbf{J}_{\ell}\in\mathbb{R}^{(\ell+1)\times(\ell+1)}$. Let $C_{i}=\frac{B_{i}}{\sqrt{1+t(1-\rho^{2})}}$ for all $0\leq i\leq\ell$ and denote $\mathbf{C}_{\ell}\triangleq[C_{0},C_{1},\cdots,C_{\ell}]$. By (47), we note that
$m_{\ell}$ $=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\frac{1}{\sqrt{1+t(1-\rho^{2})}}\exp\left(\frac{-\frac{t}{2}\left(\rho B_{i}-B_{i+1}\right)^{2}}{1+t(1-\rho^{2})}\right)\right]$
$=\int\cdots\int\frac{\sqrt{1+t(1-\rho^{2})}}{\left(\sqrt{2\pi(1+t(1-\rho^{2}))}\right)^{\ell+1}}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell}B_{i}^{2}-\frac{t}{2}\sum_{i=0}^{\ell-1}\frac{\left(\rho B_{i}-B_{i+1}\right)^{2}}{1+t(1-\rho^{2})}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}$
$=\int\cdots\int\frac{\sqrt{1+t(1-\rho^{2})}}{\left(\sqrt{2\pi}\right)^{\ell+1}}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell}(1+t(1-\rho^{2}))C_{i}^{2}-\frac{t}{2}\sum_{i=0}^{\ell-1}\left(\rho C_{i}-C_{i+1}\right)^{2}\right)\,\mathrm{d}C_{0}\cdots\mathrm{d}C_{\ell}$
$=\sqrt{1+t(1-\rho^{2})}\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\mathbf{C}_{\ell}^{\top}\mathbf{W}_{\ell}\mathbf{C}_{\ell}\right)\,\mathrm{d}C_{0}\cdots\mathrm{d}C_{\ell}$
$=\sqrt{1+t(1-\rho^{2})}\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\mathbf{C}_{\ell}^{\top}\mathbf{J}_{\ell}^{\top}\mathbf{J}_{\ell}\mathbf{C}_{\ell}\right)\,\mathrm{d}C_{0}\cdots\mathrm{d}C_{\ell}$
$=\sqrt{1+t(1-\rho^{2})}\left[\det\left(\mathbf{J}_{\ell}\right)\right]^{-1}=\left[\frac{\det\left(\mathbf{W}_{\ell}\right)}{1+t(1-\rho^{2})}\right]^{-1/2}$ (48)
We then compute the explicit formula of $\det\left(\mathbf{W}_{\ell}\right)$. Let
$\mathbf{U}_{\ell}\triangleq\begin{bmatrix}1+t&-t\rho&0&\cdots&0\\ -t\rho&1+2t&-t\rho&\cdots&0\\ 0&-t\rho&1+2t&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&-t\rho\\ 0&0&\cdots 0&-t\rho&1+2t\end{bmatrix},$
where $(\mathbf{U}_{\ell})_{00}=1+t,(\mathbf{U}_{\ell})_{ii}=1+2t$ for any $1\leq i\leq\ell$ and $(\mathbf{U}_{\ell})_{ij}=-t\rho$ for any $|i-j|=1$ with $0\leq i,j\leq\ell$. Expanding the last column of $\mathbf{W}_{\ell}$ and $\mathbf{U}_{\ell}$ yields that
$\det(\mathbf{W}_{\ell})$ $=\left(1+t(2-\rho^{2})\right)\det(\mathbf{U}_{\ell-1})-t^{2}\rho^{2}\det(\mathbf{U}_{\ell-2}),\text{ for any }\ell\geq 2,$
$\det(\mathbf{U}_{\ell})$ $=\left(1+2t\right)\det(\mathbf{U}_{\ell-1})-t^{2}\rho^{2}\det(\mathbf{U}_{\ell-2}),\text{ for any }\ell\geq 2.$
Therefore, the general form of $\det(\mathbf{U}_{\ell})$ is determined by $\det(\mathbf{U}_{\ell})=\alpha_{1}^{\prime}\lambda_{1}^{\ell}+\alpha_{2}^{\prime}\lambda_{2}^{\ell}$, where $\lambda_{1}>\lambda_{2}$ are two roots of the equation $x^{2}-(1+2t)x+t^{2}\rho^{2}=0$. Since $\lambda_{1}+\lambda_{2}=1+2t>0$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}>0$ and the discriminant $(1+2t)^{2}-4t^{2}\rho^{2}>0$, we obtain $\lambda_{1}>\lambda_{2}>0$. Consequently, the general form of $\det(\mathbf{W}_{\ell})$ is given by
$\det(\mathbf{W}_{\ell})=\alpha_{1}^{\prime\prime}\lambda_{1}^{\ell}+\alpha_{2}^{\prime\prime}\lambda_{2}^{\ell}.$
The coefficients $\alpha_{1}^{\prime\prime}$ and $\alpha_{2}^{\prime\prime}$ can be determined via the first two terms $\det(\mathbf{W}_{1})=(1+t(1-\rho^{2}))(1+2t)$ and $\det(\mathbf{W}_{2})=(1+t(1-\rho^{2}))(1+4t+t^{2}(4-\rho^{2}))$. Then we get
$\det(\mathbf{W}_{\ell})=(1+t(1-\rho^{2}))\left[\left(\frac{1}{2}+\frac{1+2t}{2\sqrt{(1+2t)^{2}-4t^{2}\rho^{2}}}\right)\lambda_{1}^{\ell}+\left(\frac{1}{2}-\frac{1+2t}{2\sqrt{(1+2t)^{2}-4t^{2}\rho^{2}}}\right)\lambda_{2}^{\ell}\right].$
Denote $\det(\mathbf{W}_{\ell})=(1+t(1-\rho^{2}))\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)$. Then, $\alpha_{1}+\alpha_{2}=1$. Since $1+2t>\sqrt{(1+2t)^{2}-4t^{2}\rho^{2}}$, we obtain that $\alpha_{1}>0$ and $\alpha_{2}<0$. Then, by (48),
$m_{\ell}=\left[\frac{\det\left(\mathbf{W}_{\ell}\right)}{1+t(1-\rho^{2})}\right]^{-1/2}=\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)^{-1/2}.$
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for $i=0,\dots,\ell-1$. Recall that $\lambda_{1}+\lambda_{2}=1+2t$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}$. Following a similar argument as (47), we have
$\tilde{m}_{\ell}$ $\triangleq\mathbb{E}\left[e^{t\beta_{C}(\mathcal{H}_{\pi}^{f})}\right]=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\frac{1}{\sqrt{1+t(1-\rho^{2})}}\exp\left(\frac{-\frac{t}{2}\left(\rho B_{i}-B_{i+1}\right)^{2}}{1+t(1-\rho^{2})}\right)\right]$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi(1+t(1-\rho^{2}))}}\right)^{\ell}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell-1}B_{i}^{2}-\frac{t}{2}\sum_{i=0}^{\ell-1}\frac{\left(\rho B_{i}-B_{i+1}\right)^{2}}{1+t(1-\rho^{2})}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell-1}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell-1}(1+2t)C_{i}^{2}+\frac{1}{2}\sum_{i=0}^{\ell-1}t\rho C_{i}C_{i+1}\right)\,\mathrm{d}C_{0}\cdots\mathrm{d}C_{\ell-1}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell-1}\left(\lambda_{1}^{1/2}C_{i}-\lambda_{2}^{1/2}C_{i+1}\right)^{2}\right)\,\mathrm{d}C_{0}\cdots\mathrm{d}C_{\ell-1}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell}\exp\left(-\frac{1}{2}\mathbf{C}_{\ell-1}^{\top}\tilde{\mathbf{J}}_{\ell-1}^{\top}\tilde{\mathbf{J}}_{\ell-1}\mathbf{C}_{\ell-1}\right)\,\mathrm{d}C_{0}\cdots\mathrm{d}C_{\ell-1}=\left[\det(\tilde{\mathbf{J}}_{\ell-1})\right]^{-1},$
where
$\tilde{\mathbf{J}}_{\ell-1}=\begin{bmatrix}\lambda_{1}^{1/2}&-\lambda_{2}^{1/2}&0&\cdots&0\\ 0&\lambda_{1}^{1/2}&-\lambda_{2}^{1/2}&\cdots&0\\ 0&0&\lambda_{1}^{1/2}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&-\lambda_{2}^{1/2}\\ -\lambda_{2}^{1/2}&0&\cdots 0&0&\lambda_{1}^{1/2}\end{bmatrix}$
and hence $\det(\tilde{\mathbf{J}}_{\ell-1})=\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}$. Therefore,
$m_{\ell}^{-2}-\tilde{m}_{\ell}^{-2}$ $=\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}-\left(\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}\right)^{2}$
$=(\alpha_{1}-1)\left(\lambda_{1}^{\ell}-\lambda_{2}^{\ell}\right)+\lambda_{2}^{\ell/2}\left(\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}\right)+\lambda_{1}^{\ell/2}\lambda_{2}^{\ell/2}>0,$
which implies that $m_{\ell}<\tilde{m}_{\ell}$ for any $\ell\geq 1$.
Next, we prove that $\tilde{m}_{\ell}\leq\left(\tilde{m}_{2}\right)^{\ell/2}$ for $\ell\geq 2$. Indeed, $\tilde{m}_{\ell}^{-1}=\left(\lambda_{2}+(\lambda_{1}-\lambda_{2})\right)^{\ell/2}-\lambda_{2}^{\ell/2}\geq(\lambda_{1}-\lambda_{2})^{\ell/2}=\tilde{m}_{2}^{-\ell/2}$, and $\tilde{m}_{2}^{-1}-\tilde{m}_{1}^{-2}=2\lambda_{2}^{1/2}\left(\lambda_{1}^{1/2}-\lambda_{2}^{1/2}\right)>0$. Recall that $\kappa_{\ell}^{\mathsf{C}}(t)=\log\left(\tilde{m}_{\ell}\right),\kappa_{\ell}^{\mathsf{C}}(t)=\log\left(m_{\ell}\right)$, and $m_{\ell}<\tilde{m}_{\ell}$. Consequently,
$\frac{1}{2}\kappa_{2}^{\mathsf{C}}(t)\leq\kappa_{1}^{\mathsf{C}}(t),\quad\kappa_{\ell}^{{\mathsf{P}}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)\quad\forall\ell\geq 2.$
The last inequality $\kappa_{1}^{\mathsf{P}}(t)\leq\frac{1}{2}\kappa_{2}^{\mathsf{C}}(t)$ follows from
$m_{1}^{-4}-\tilde{m}_{2}^{-2}=\left(1+2t\right)^{2}-\left[(1+2t)^{2}-4t^{2}\rho^{2}\right]=4t^{2}\rho^{2}>0.$
Finally, since the summands over different connected components are independent, it follows that
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})}\right]$ $=\sum_{P\in{\mathcal{P}}}\kappa_{|P|}^{{\mathsf{P}}}(t)+\sum_{C\in{\mathcal{C}}}\kappa_{|C|}^{{\mathsf{C}}}(t)$
$\leq\sum_{P\in{\mathcal{P}}}\frac{|P|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+\sum_{C\in{\mathcal{C}}:|C|\geq 2}\frac{|C|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+\sum_{C\in{\mathcal{C}}:|C|=1}\kappa_{1}^{{\mathsf{C}}}(t)$
$=\frac{|{\mathcal{E}}|}{2}\kappa_{2}^{{\mathsf{C}}}(t)+|\{C\in{\mathcal{C}}:|C|=1\}|\left(\kappa_{1}^{{\mathsf{C}}}(t)-\frac{1}{2}\kappa_{2}^{{\mathsf{C}}}(t)\right),$
where the last equality uses fact that $|{\mathcal{E}}|=\sum_{P\in{\mathcal{P}}}|P|+\sum_{C\in{\mathcal{C}}}|C|$.
Remark 6.
We have two bounds for large $\ell$ in the Gaussian Wigner model, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)\geq\frac{1}{\ell}\log\left(\left(\lambda_{1}^{\ell/2}-\lambda_{2}^{\ell/2}\right)^{2}\right)$, where $\alpha_{2}<0<1<\alpha_{1},\alpha_{1}+\alpha_{2}=1$ and $\lambda_{1}>\lambda_{2}>0$. Consequently, $-2\left(\lambda_{1}+\frac{\log\alpha_{1}}{\ell}\right)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq-2\left(\lambda_{1}+\frac{1}{\ell}\log\left[1-2\left(\frac{\lambda_{2}}{\lambda_{1}}\right)^{\ell/2}\right]\right).$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $(x+y)^{\ell/2}-y^{\ell/2}\geq x^{\ell/2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
B.2 Proof of Lemma 2
We first upper bound the higher-order cumulants. By Lemma 1, for any $t>0$,
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})}\right]\leq\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$
where $L$ denotes the number of self-loops. The self-loop for $e$ only happens when $\pi(e)=\pi^{*}(e)$. For $uv\in\binom{V(G_{1})}{2}\backslash\binom{F_{\pi}}{2}$, by the definition of $F_{\pi}$, $\pi(u)\neq\pi^{*}(u)$ or $\pi(v)\neq\pi^{*}(v)$. Therefore, $\pi(uv)=\pi^{*}(uv)$ implies that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$. Since $d(\pi^{*},\pi)=k$, we must have $L\leq\frac{k}{2}$. Applying the formulas (36) and (37) and the fact that $p_{11}\leq p$, we obtain
$\kappa^{{\mathsf{C}}}_{2}(t)$ $\leq\log\left(1+2p^{2}(e^{t}-1)+p^{2}(e^{t}-1)^{2}\right)=\log\left(1+p^{2}(e^{2t}-1)\right)\leq p^{2}(e^{2t}-1)$
and
$\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)$ $=\frac{1}{2}\log\left[1+\frac{2(p_{11}-p^{2})}{p_{11}^{2}(e^{t}-1)+2p^{2}+(e^{t}-1)^{-1}}\right]$
$\overset{\mathrm{(a)}}{\leq}\frac{1}{2}\log\left[1+\frac{2(p_{11}-p^{2})}{2(p_{11}+p^{2})}\right]\overset{\mathrm{(b)}}{\leq}\frac{\gamma}{2(\gamma+2)},$
where $\mathrm{(a)}$ is because $p_{11}^{2}(e^{t}-1)+(e^{t}-1)^{-1}\geq 2\sqrt{p_{11}^{2}(e^{t}-1)(e^{t}-1)^{-1}}=2p_{11}$ and $\mathrm{(b)}$ is because $p_{11}=p^{2}(1+\gamma)$ and $\log(1+x)\leq x$ for any $x\geq 0$. Therefore, we obtain
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})}\right]\leq\frac{|{\mathcal{E}}_{\pi}|}{2}p^{2}(e^{2t}-1)+\frac{k\gamma}{4(\gamma+2)}.$
We then apply the Chernoff bound the provide an upper bound for $\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]$. For any $t>0$,
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-t\tau_{k}+\frac{|{\mathcal{E}}_{\pi}|}{2}p^{2}(e^{2t}-1)+\frac{k\gamma}{4(2+\gamma)}\right).$
Let $t=\frac{1}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)$. Then $t>0$ by the assumption $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$. We obtain that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}+\frac{k\gamma}{4(2+\gamma)}\right).$
B.3 Proof of Lemma 3
We first upper bound the higher-order cumulants. Let $t=\frac{\rho}{3\sqrt{1+\rho^{2}}}$. Then $0<t<\frac{1}{1+\rho}$ since $10<\rho<$ 1. By Lemma 1,
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})}\right]\leq\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$
where $L$ denotes the number of self-loops. The self-loop for $e$ only happens when $\pi(e)=\pi^{*}(e)$. For $uv\in\binom{V(G_{1})}{2}\backslash\binom{F_{\pi}}{2}$, by the definition of $F_{\pi}$, $\pi(u)\neq\pi^{*}(u)$ or $\pi(v)\neq\pi^{*}(v)$. Therefore, $\pi(uv)=\pi^{*}(uv)$ implies that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$. Since $d(\pi^{*},\pi)=k$, we must have $L\leq\frac{k}{2}$. Applying the formulas (41) and (42), we obtain that
$\kappa_{1}^{\mathsf{C}}(t)-\frac{1}{2}\kappa_{2}^{\mathsf{C}}(t)=$ $\frac{1}{4}\log\left[\frac{\left(1-t^{2}(1-\rho^{2})-2t\rho\right)\left(1-t^{2}(1-\rho^{2})+2t\rho\right)}{\left(1-t^{2}(1-\rho^{2})-2t\rho\right)^{2}}\right]$
$=$ $\frac{1}{4}\log\left[1+\frac{4t\rho}{1-t^{2}(1-\rho^{2})-2t\rho}\right]\overset{\mathrm{(a)}}{\leq}\frac{1}{4}\log\left(1+\frac{12\rho^{2}}{9-7\rho^{2}+\rho^{4}}\right)\overset{\mathrm{(b)}}{\leq}\frac{\log 5}{4},$
where $\mathrm{(a)}$ is because $\frac{4t\rho}{1-t^{2}(1-\rho^{2})-2t\rho}=\frac{12\rho^{2}}{9\sqrt{1+\rho^{2}}-\frac{\rho^{2}(1-\rho^{2})}{\sqrt{1+\rho^{2}}}-6\rho^{2}}\leq\frac{12\rho^{2}}{9-\rho^{2}(1-\rho^{2})-6\rho^{2}}=\frac{12\rho^{2}}{9-7\rho^{2}+\rho^{4}}$; $\mathrm{(b)}$ is because $\frac{12x^{2}}{9-7x^{2}+x^{4}}$ is increasing on $(0,1)$ and $\rho<1$.
By the Chernoff bound, we have that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)\geq\tau_{k}\right]\leq\exp\left(-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right)\right).$
It remains to upper bound $-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$. We note that
$-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$ $=-t\tau_{k}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left[1-\frac{2\rho^{2}}{9}+\frac{\rho^{4}(1-\rho^{2})^{2}}{81(1+\rho^{2})^{2}}\right]$
$\overset{\mathrm{(a)}}{\leq}-t\tau_{k}+\frac{9|\mathcal{E}_{\pi}|}{28}\left[\frac{2\rho^{2}}{9}-\frac{\rho^{4}(1-\rho^{2})^{2}}{81(1+\rho^{2})^{2}}\right]\overset{\mathrm{(b)}}{\leq}-\frac{\rho\tau_{k}}{6}+\frac{\rho^{2}|\mathcal{E}_{\pi}|}{14},$
where $\mathrm{(a)}$ is because $\frac{2\rho^{2}}{9}-\frac{\rho^{4}(1-\rho^{2})^{2}}{81(1+\rho^{2})^{2}}=\frac{\rho^{2}\left(18(1+\rho^{2})^{2}-\rho^{2}\left(1-\rho^{2}\right)^{2}\right)}{81(1+\rho^{2})^{2}}\geq 0$ and $\log(1-x)\geq-\frac{9x}{7}$ for any $0\leq x\leq\frac{2}{9}$; $\mathrm{(b)}$ is because $t\geq\frac{\rho}{6}$. Therefore, we obtain that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)\geq\tau_{k}\right]\leq\exp\left(-\frac{\rho\tau_{k}}{6}+\frac{\rho^{2}|\mathcal{E}_{\pi}|}{14}+\frac{\log 5}{8}k\right).$
B.4 Proof of Lemma 4
We first upper bound the higher-order cumulants. By Lemma 1, for any $t>0$,
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)}\right]\leq\frac{|{\mathcal{E}}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$
where $L$ denotes the number of self-loops. The self-loop for $e$ only happens when $\pi(e)=\pi^{*}(e)$. For $uv\in\binom{V(G_{1})}{2}\backslash\binom{F_{\pi}}{2}$, by the definition of $F_{\pi}$, $\pi(u)\neq\pi^{*}(u)$ or $\pi(v)\neq\pi^{*}(v)$. Therefore, $\pi(uv)=\pi^{*}(uv)$ implies that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$. Since $d(\pi^{*},\pi)=k$, we must have $L\leq\frac{k}{2}$. Let $t=\frac{\rho}{4(1-\rho)}$. Applying the formulas (45) and (46), we obtain that
$\kappa_{1}^{\mathsf{C}}(t)-\frac{1}{2}\kappa_{2}^{\mathsf{C}}(t)$ $=\frac{1}{4}\log\left[\frac{(1+2t-2t\rho)(1+2t+2t\rho)}{(1+2t(1-\rho))^{2}}\right]$
$=\frac{1}{4}\log\left(1+\frac{4t\rho}{1+2t-2t\rho}\right)\leq\frac{1}{4}\log\left(\frac{1}{1-\rho}\right),$
where the last inequality follows from $1+\frac{4t\rho}{1+2t-2t\rho}=\frac{2+\rho^{2}-\rho}{2+\rho}\cdot\frac{1}{1-\rho}\leq\frac{1}{1-\rho}$.
By the Chernoff bound, we have that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)\geq\tau_{k}\right]\leq\exp\left(-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right)\right).$
It remains to upper bound $-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$. We note that
$-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$ $=-t\tau_{k}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left[(1+2t)^{2}-4t^{2}\rho^{2}\right]$
$\leq-t\tau_{k}-\frac{|\mathcal{E}_{\pi}|}{4}\log(1+4t)=-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right).$
Therefore, we obtain that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right)+\frac{k}{8}\log\left(\frac{1}{1-\rho}\right)\right).$
Appendix C Auxiliary results
Lemma 5.
Recall that $\phi(\gamma)=(1+\gamma)\log(1+\gamma)-\gamma$ and $\eta,\gamma>0$. If $\eta\leq\frac{\gamma}{4(1+\gamma)}$, then $(1+\gamma)(1-\eta)>1$ and
$\phi\left[(1-\eta)(1+\gamma)-1\right]\geq\frac{1}{4}\phi(\gamma).$ (49)
Proof.
We note that $(1+\gamma)(1-\eta)\geq 1+\gamma-\frac{\gamma}{4}>1$, and
$\phi\left[(1-\eta)(1+\gamma)-1\right]$ $=(1+\gamma)(1-\eta)\log\left[(1+\gamma)(1-\eta)\right]-\left[(1+\gamma)(1-\eta)-1\right]$
$=(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+(1+\gamma)(1-\eta)\log(1-\eta)+\eta$
$\geq(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+(1+\gamma)(-\eta)+\eta$
$=(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]-\eta\gamma,$
where the last inequality is due to the fact that $(1-x)\log(1-x)+x\geq 0$ for any $10<x<$ 1 and $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. Since $\eta\leq\frac{\gamma}{4(1+\gamma)}$ and $(1+\gamma)\log(1+\gamma)-\gamma\geq\frac{\gamma^{2}}{2(1+\gamma)}$, we obtain $\eta\gamma\leq\frac{\gamma^{2}}{4(1+\gamma)}\leq\frac{1}{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$. Therefore,
$\phi\left[(1-\eta)(1+\gamma)-1\right]$ $\geq(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]-\eta\gamma$
$\geq\left(\frac{1}{2}-\eta\right)\phi(\gamma)\geq\frac{1}{4}\phi(\gamma),$
where the last inequality is because $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. ∎
Lemma 6.
For any $m\geq 10$,
$\sum_{k=1}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]\leq\frac{4\log m+2}{m},$
where $h(x)=-x\log x-(1-x)\log(1-x)$ is the binary entropy function.
Proof.
We note that
$\sum_{k=1}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]$ $\overset{\mathrm{(a)}}{\leq}2\sum_{1\leq k\leq\frac{m}{2}}\exp\left[-k\log\left(\frac{m}{k}\right)\right]$
$\overset{\mathrm{(b)}}{\leq}2\sum_{1\leq k\leq 2\log m}\exp\left[-k\log\left(\frac{m}{k}\right)\right]+2\sum_{2\log m+1\leq k\leq\frac{m}{2}}2^{-k}$
$\overset{\mathrm{(c)}}{\leq}2\cdot\exp\left(-\log m\right)\cdot(2\log m)+2\cdot 2^{-2\log m}$
$\overset{\mathrm{(d)}}{\leq}\frac{4\log m+2}{m},$
where $\mathrm{(a)}$ is because $h(x)=h(1-x)$ and $h(x)\geq-x\log x$; $\mathrm{(b)}$ is because $\log\left(\frac{m}{k}\right)\geq\log 2$ when $k\leq\frac{m}{2}$; $\mathrm{(c)}$ is because $k\log\left(\frac{m}{k}\right)\geq\log m$ for $1\leq k\leq 2\log m$ when $m\geq 10$; $\mathrm{(d)}$ is because $2\cdot 2^{-2\log m}\leq\frac{2}{m}$. ∎
Lemma 7.
For $P_{1}=\mathrm{Bern}(p,p,\rho)$ and $Q_{1}=\mathrm{Bern}(p,p,0)$, we have
$D(P_{1}\|Q_{1})\leq 25p^{2}\phi(\gamma).$
For $P_{2}=\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ and $Q_{2}=\mathcal{N}\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right)$, we have
$D(P_{2}\|Q_{2})=\frac{1}{2}\log\left(\frac{1}{1-\rho^{2}}\right).$
Proof.
Recall $p_{ab}$ for any $a,b\in\{0,1\}$ defined in (4). Then, we have
$D(P_{1}\|Q_{1})=$$~{}\sum_{\{a,b\}\in\{0,1\}}p_{ab}\log\left[\frac{p_{ab}}{p^{a+b}(1-p)^{2-a-b}}\right]$
$=$ $~{}\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]+2p(1-p)(1-\rho)\log(1-\rho)$
$~{}+\left[(1-p)^{2}+\rho p(1-p)\right]\log\left(1+\frac{\rho p}{1-p}\right)$
$\overset{\mathrm{(a)}}{\leq}$ $~{}\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]+2p(1-p)(1-\rho)\cdot(-\rho)$
$~{}+\left[(1-p)^{2}+\rho p(1-p)\right]\cdot\frac{\rho p}{1-p}$
=$\overset{\mathrm{(b)}}{=}$ $~{}p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right],$
where $\mathrm{(a)}$ uses $\log(1+x)\leq x$ for any $x>-1$; $\mathrm{(b)}$ follows from $\gamma=\frac{\rho(1-p)}{p}$. Since $\log(1+x)\geq\frac{x}{x+1}+\frac{x^{2}}{2(x+1)^{2}}$ for any $x\geq 0$, we obtain that $p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\geq\frac{p^{2}\gamma^{2}}{2(\gamma+1)}$. When $\gamma<3$, since $\frac{p^{2}\gamma^{2}}{2(\gamma+1)}\geq\frac{p^{2}\gamma^{2}}{8}=\frac{\rho^{2}\cdot 3(1-p)^{2}}{24}\geq\frac{\rho^{2}\left[2p(1-p)+p^{2}\right]}{24}$ for $0<p\leq\frac{1}{2}$, we obtain that
$D(P_{1}\|Q_{1})\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right]\leq 25p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right].$
When $\gamma\geq 3$, since $\rho^{2}\left[2p(1-p)+p^{2}\right]\leq 3\rho p(1-p)$ and $(\log 4-1)\rho p(1-p)=p^{2}\gamma(\log 4-1)\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$, we obtain
$D(P_{1}\|Q_{1})$ $\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right]$
$\leq\left(\frac{3}{\log 4-1}+1\right)p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$
$\leq 25p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right].$
Therefore, we get $D(P_{1}\|Q_{1})\leq 25p^{2}\phi(\gamma)$.
We denote $P_{2}(a,b)$ and $Q_{2}(a,b)$ the probability density function under $P_{2}$ and $Q_{2}$, respectively. Then, the KL-divergence between $P_{2}$ and $Q_{2}$ is given by
$D(P_{2}\|Q_{2})$ $=\iint P_{2}(a,b)\log\left(\frac{P_{2}(a,b)}{Q_{2}(a,b)}\right)\,\mathrm{d}a\mathrm{d}b$
$=\iint P_{2}(a,b)\left[\frac{1}{2}\log\left(\frac{1}{1-\rho^{2}}\right)+\frac{\rho ab}{1-\rho^{2}}-\frac{\rho^{2}(a^{2}+b^{2})}{2(1-\rho^{2})}\right]\,\mathrm{d}a\mathrm{d}b$
$=\frac{1}{2}\log\left(\frac{1}{1-\rho^{2}}\right)+\frac{\rho^{2}}{1-\rho^{2}}-\frac{2\rho^{2}}{2(1-\rho^{2})}=\frac{1}{2}\log\left(\frac{1}{1-\rho^{2}}\right).\qed$
Lemma 8 (Chernoff’s inequality for Binomials).
Suppose $\xi\sim\mathrm{Bin}(n,p)$, denote $\mu=np$, for any $\delta>0$,
$\mathbb{P}\left[\xi\geq(1+\delta)\mu\right]$ $\leq\exp\left\{-\mu\left[(1+\delta)\log(1+\delta)-\delta\right]\right\},$ (50)
$\mathbb{P}\left[\xi\geq(1+\delta)\mu\right]$ $\leq\exp\left(-\frac{\delta\mu}{2+\delta}\right).$ (51)
For any $10<\delta<$ 1, we have
$\mathbb{P}\left[\xi\leq(1-\delta)\mu\right]\leq\exp\left(-\frac{\delta^{2}\mu}{2}\right).$ (52)
Proof.
By Theorems 4.4 and 4.5 in [MU05] we have (50) and (52). Since $(1+\delta)\log(1+\delta)-\delta\geq\frac{\delta^{2}}{2+\delta}$, we obtain (51) from (50). ∎
Lemma 9 (Hanson-Wright inequality).
Let $X,Y\in\mathbb{R}^{n}$ be standard Gaussian vectors such that the pairs $(X_{i},Y_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ are independent for $i=1,\cdots,n$. Let $M_{0}\in\mathbb{R}^{n\times n}$ be any deterministic matrix. There exists some universal constant $c_{0}>0$ such that,
$\mathbb{P}\left[\left|X^{\top}M_{0}Y-\rho\mathrm{Tr}(M_{0})\right|\geq c_{0}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]\leq\delta.$
Proof.
Note that $X^{\top}M_{0}Y=\frac{1}{4}(X+Y)^{\top}M_{0}(X+Y)-\frac{1}{4}(X-Y)^{\top}M_{0}(X-Y)$ and
$\mathbb{E}\left[(X+Y)^{\top}M_{0}(X+Y)\right]=(2+2\rho)\mathrm{Tr}(M_{0}),\mathbb{E}\left[(X-Y)^{\top}M_{0}(X-Y)\right]=(2-2\rho)\mathrm{Tr}(M_{0}).$
By Hanson-Wright inequality [HW71], there exists some universal constant $c_{0}$ such that
$\mathbb{P}\left[\left|\frac{1}{4}(X+Y)^{\top}M_{0}(X+Y)-\frac{2+2\rho}{4}\mathrm{Tr}(M_{0})\right|\geq\frac{c_{0}}{2}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]\leq\frac{\delta}{2},$
$\mathbb{P}\left[\left|\frac{1}{4}(X-Y)^{\top}M_{0}(X-Y)-\frac{2-2\rho}{4}\mathrm{Tr}(M_{0})\right|\geq\frac{c_{0}}{2}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]\leq\frac{\delta}{2}$
for any $\delta>0$. Consequently,
$~{}\mathbb{P}\left[\left|X^{\top}M_{0}Y-\rho\mathrm{Tr}(M_{0})\right|\geq c_{0}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]$
$\leq$ $~{}\mathbb{P}\left[\left|\frac{1}{4}(X+Y)^{\top}M_{0}(X+Y)-\frac{2+2\rho}{4}\mathrm{Tr}(M_{0})\right|\geq\frac{c_{0}}{2}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]$
$+\mathbb{P}\left[\left|\frac{1}{4}(X-Y)^{\top}M_{0}(X-Y)-\frac{2-2\rho}{4}\mathrm{Tr}(M_{0})\right|\geq\frac{c_{0}}{2}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]$
$\leq$ $~{}\delta.\qed$
Lemma 10 (Chernoff’s inequality for Chi-squared distribution).
Suppose $\xi$ follows the chi-squared distribution with $n$ degrees of freedom. Then, for any $\delta>0$,
$\mathbb{P}\left[\xi>(1+\delta)n\right]\leq\exp\left(-\frac{n}{2}\left(\delta-\log\left(1+\delta\right)\right)\right).$ (53)
Proof.
The result follows from [Gho21, Theorem 1]. ∎